### v1

# CodeWiki: Automated Repository-Level Documentation at Scale

Abstract
Developers spend nearly 58% of their time understanding codebases, yet maintaining comprehensive documentation remains challenging due to complexity and manual effort. While recent Large Language Models (LLMs) show promise for function-level documentation, they fail at the repository level, where capturing architectural patterns and cross-module interactions is essential. We introduce CodeWiki, the first open-source framework for holistic repository-level documentation across seven programming languages. CodeWiki employs three innovations: (i) hierarchical decomposition that preserves architectural context, (ii) recursive agentic processing with dynamic delegation, and (iii) synthesis of textual and visual artifacts including architecture diagrams and data flows. We also present CodeWikiBench, the first repository-level documentation benchmark with multi-level rubrics and agentic assessment. CodeWiki achieves 68.79% quality score with proprietary models and 64.80% with open-source alternatives, outperforming existing closed-source systems and demonstrating scalable, accurate documentation for real-world repositories.
1 Introduction
Figure 1: CodeWiki Framework Architecture Overview. The framework operates in three main phases: (1) Repository analysis through AST/LLM parsing to construct dependency graphs and identify high-level components, followed by hierarchical decomposition into manageable modules; (2) Recursive documentation generation where specialized agents process leaf modules with dynamic delegation capabilities, creating markdown documentation while maintaining cross-module references; (3) Hierarchical assembly where parent modules are synthesized from child documentation using LLM-based synthesis, culminating in comprehensive repository overview documentation. The module tree evolves throughout the process, enabling scalable processing of repositories of arbitrary size.
In the rapidly evolving landscape of software development, maintaining comprehensive and up-to-date documentation has become increasingly challenging as codebases grow in size and complexity. Studies indicate that developers spend approximately 58% of their working time on program comprehension activities (Xia et al., 2018), highlighting the critical importance of accessible and accurate documentation. Despite documentation being recognized as essential for software maintenance and collaboration (de Souza et al., 2005; Zhi et al., 2015; Chen and Huang, 2009), the manual creation and maintenance of repository-level documentation remain a labor-intensive and often neglected aspect of software engineering practice (McBurney et al., 2018). The emergence of Large Language Models (LLMs)  (Zhang et al., 2023; He et al., 2025b) has opened opportunities for automating documentation generation tasks. However, current approaches still face two fundamental limitations.
Limitations of Scalability at the Repository Level
Recent advances in code understanding and natural language generation have demonstrated promising results in function-level and file-level documentation tasks (Feng et al., 2020; Khan and Uddin, 2023; Poudel et al., 2024; Makharev and Ivanov, 2025; Lomshakov et al., 2024; Ahmed and Devanbu, 2023; Choi et al., 2023; Luo et al., 2024; Yang et al., 2025) However, existing approaches face significant challenges when scaling to repository-level documentation.
Unlike function-level documentation that focuses on individual components, repository-level documentation task is to capture architectural patterns, cross-module interactions, data flows, and system-wide design decisions (Rai et al., 2022; Treude et al., 2020). Furthermore, the hierarchical nature of software systems requires documentation that can serve multiple audiences, from high-level stakeholders seeking architectural overviews to developers requiring detailed implementation guidance (Nassif and Robillard, 2025).
$\hookrightarrow$ To address these challenges, we present CodeWiki, the first open-source framework specifically designed to generate holistic, structured repository-level documentation that includes diverse content types such as system architecture diagrams, data flow visualizations, and sequence diagrams while scaling effectively for projects of arbitrary size. CodeWiki introduces several key innovations: a hierarchical decomposition strategy inspired by dynamic programming principles that breaks complex repositories into manageable modules while preserving architectural coherence, a recursive agentic system with dynamic delegation capabilities that enables adaptive processing based on module complexity, and a comprehensive documentation synthesis approach that generates multiple forms of visual and textual documentation reflecting fine-grained codebase collaboration patterns.
Limitations of the Evaluation Landscape
Despite progress, systematic benchmarks remain lacking. Function-level and file-level work relies on general-purpose metrics (BLEU, ROUGE) that fail to capture documentation quality nuances (Roy et al., 2021; Evtikhiev et al., 2023). Repository-level evaluation is particularly challenging as documentation admits multiple valid structures sharing underlying architectural content, highlighting the need for robust assessment beyond surface-level similarity.
$\hookrightarrow$ To fill this gap, we introduce CodeWikiBench, the first benchmark specifically designed for repository-level documentation. Instead of relying on surface-level similarity, we derive hierarchical rubrics from the official documentation maintained by project developers, enabling the evaluation of underlying repository properties. For robustness, CodeWikiBench further adopts an agentic assessment framework in which judge agents evaluate generated documentation against rubric requirements and aggregate results hierarchically, producing both quality scores and reliability measures.
Limitations of Multilingual Support
Most studies on code documentation generation have primarily focused on Python, with limited consideration of other widely used programming languages such as Java, JavaScript, C, or C++. This narrow scope restricts generalizability and overlooks the structural diversity inherent in real-world software projects, where repositories often involve multiple languages.
$\hookrightarrow$ CodeWiki supports repository-level documentation generation across 7 programming languages (Python, Java, JavaScript, TypeScript, C, C++ and C#). In addition, CodeWikiBench provides a multilingual benchmark built from multiple repositories in the same set of languages, enabling comprehensive evaluation.
Using CodeWikiBench, we demonstrate CodeWiki achieves 68.79% quality score with proprietary models, outperforming closed-source DeepWiki, while open-source alternatives achieve 64.80%, producing comprehensive documentation with architectural diagrams, usage patterns, and dependency visualizations (Appendix A).
In summary, the main contributions of this work are:
  1. 1.
    CodeWiki, the first open-source semi-agentic framework for scalable repository-level documentation with proprietary and open-source LLM backends.
  2. 2.
    CodeWikiBench, the first repository-level documentation benchmark with hierarchical rubrics and agentic assessment.
  3. 3.
    First multilingual repository documentation supporting 7 widely used languages.
  4. 4.
    Competitive performance on real-world repositories, improving over DeepWiki with both proprietary (Claude Sonnet 4) and open-source (Kimi K2 Instruct) models.
2 Our Approach - CodeWiki
CodeWiki is a semi-agentic framework that automatically generates comprehensive repository-level documentation by addressing context limitations through hierarchical decomposition. As illustrated in Figure 1, our approach operates through three phases: repository analysis and module decomposition, recursive documentation generation, and hierarchical assembly, inspired by dynamic programming principles.
2.1 Repository Analysis and Hierarchical Module Decomposition
We begin with static analysis to construct a dependency graph capturing structural relationships, crucial for understanding architecture and identifying decomposition boundaries.
Dependency Graph Construction
Following DocAgent (Yang et al., 2025), we employ Tree-Sitter parsers to extract ASTs, systematically identifying functions, methods, classes, structs, modules, and interdependencies: function calls, class inheritance, attribute access, and module imports. We normalize these to a unified depends_on relation for cross-language generalization, creating directed graph $G=(V,E)$ where edge $(u,v)\in E$ indicates component $u$ depends on $v$ (Liu et al., 2023).
Entry Point Identification and Decomposition
Topological sorting identifies zero-in-degree components—independent entry points where users interact (main functions, API endpoints, CLI, public interfaces). High-level components are hierarchically decomposed into manageable modules through recursive partitioning considering component interdependencies and semantic coherence. For scalability, only component IDs serve as input, resulting in a feature-oriented module tree.
2.2 Recursive Documentation Generation
CodeWiki’s core innovation is recursive agent-based processing, enabling arbitrary repository size handling while maintaining bounded complexity and architectural coherence.
Agent Architecture
Each leaf module is assigned a specialized agent equipped with: (1) complete source code access, (2) full module tree for cross-module understanding, (3) documentation workspace tools (view, create, edit operations), and (4) dependency graph traversal for contextual exploration.
The agent workflow: analyzes components to understand functionality and interfaces, explores context by traversing dependencies within and across modules, and generates comprehensive markdown documentation including descriptions, usage examples, API specifications, and architectural insights.
Dynamic Delegation
Input: Module tree $T$, dependency graph $G$
Output: Complete repository documentation
Initialize documentation workspace;
for each leaf module $m$ in $T$ do
   agent $\leftarrow$ CreateAgent($m$, $G$, $T$);
 while $m$ not fully processed do
      agent.GenerateDoc($m$.name);
    if delegation requested then
         sub $\leftarrow$ agent.GetDelegationSpecs();
         UpdateModuleTree($T$, $m$, sub);
       foreach new leaf $m^{\prime}$ in sub do
            Recursively process $m^{\prime}$;
        ReviseDoc($m$.name);
ProcessParentModules($T$);
return Repository documentation;
Algorithm 1 Recursive Module Documentation
Our key innovation is adaptive scalability through dynamic delegation. When module complexity exceeds single-pass capacity, agents delegate sub-modules to specialized sub-agents. Delegation criteria: code complexity metrics (cyclomatic complexity, nesting depth), semantic diversity (functionally distinct subcomponents), and context window utilization.
The recursive process follows bottom-up processing (Algorithm 1). Upon delegation, agents provide sub-module specifications, the module tree updates, and newly created leaves process recursively. This handles modules of any size while maintaining quality.
Cross-Module Reference Management
Maintaining coherence across boundaries requires sophisticated reference management. When agents encounter external components during traversal, our intelligent resolution system creates cross-references rather than duplicating content. A global registry tracks documented components and locations, enabling quick identification of documented components.
This ensures conciseness while preserving completeness, creating interconnected documentation reflecting the actual codebase structure, enabling easy navigation between related components.
2.3 Hierarchical Assembly and Documentation Synthesis
After leaf documentation, hierarchical assembly synthesizes component-level details into architectural overviews through recursive parent module processing.
Parent modules undergo LLM synthesis with carefully crafted prompting. LLMs receive: child module documentation, module tree structure, dependency information, and synthesis instructions for architectural patterns and feature interactions.
Synthesis involves multiple stages: analyzing child documentation for themes and patterns, generating architectural overviews explaining module collaboration, creating feature summaries distilling capabilities, developing usage guides for public interfaces, and producing architectural diagrams visualizing relationships and data flows.
3 The First Benchmarking System - CodeWikiBench
Figure 2: Example hierarchical evaluation rubric for the RAGFlow repository. The rubric mirrors the project’s architectural structure with weighted requirements at multiple levels. Leaf nodes represent specific requirements assessed by Judge Agent, while parent scores are computed through weighted aggregation. The hierarchical organization ensures comprehensive coverage from high-level architectural components down to specific implementation details, with weights reflecting the relative importance of each component for understanding the repository
Repository-level documentation evaluation presents unique challenges due to complexity and hierarchical organization. Unlike code generation assessable through automated metrics, documentation quality requires nuanced evaluation of completeness, accuracy, and coherence across abstraction levels. Recent research highlights traditional metrics’ limitations for code tasks (Evtikhiev et al., 2023; Haldar and Hockenmaier, 2024), while LLM-as-a-Judge paradigms offer new possibilities (Wang et al., 2025; Crupi et al., 2025; He et al., 2025a). We develop a comprehensive framework synthesizing repository-specific rubrics and employing agentic assessment.
3.1 Hierarchical Rubric Generation
Following (Starace et al., 2025), we synthesize evaluation rubrics tailored for each repository, illustrated in Figure 2. We collect official documentation from selected open-source projects and hierarchically parse into structured JSON format, leveraging directory structure and markdown syntax. This captures multi-level technical documentation from architectural overviews to implementation details.
A Rubric Generator Agent processes each structure with tool access to detailed contents, generating hierarchical rubrics mirroring repository functionality and aligning with actual documentation standards. To enhance reliability, we employ multiple independent generations using different model families. Final rubrics emerge from synthesizing these perspectives, reducing single-model biases. Analysis shows generated rubrics demonstrate high consistency: 73.65% semantic reliability and 70.84% structural reliability (Appendix C).
3.2 Documentation Assessment
Our evaluation employs specialized Judge Agents operating under carefully designed protocols. Judges receive generated documentation in structured JSON format with hidden detailed contents, plus repository-specific rubrics. The systematic methodology: analyzes documentation structure and content, searches comprehensively for requirement coverage, makes binary adequacy decisions, and provides concise reasoning.
Critically, judges evaluate only leaf-level requirements, ensuring assessment grounds in specific concrete criteria rather than abstract concepts. For example, in Figure 2, agents assess "DeepDoc Visual Parser for complex layouts" rather than broader "Document Processing Engine."
For enhanced reliability, multiple Judge Agents using different model families evaluate each requirement. Final leaf scores average across assessments, reducing individual model biases.
3.3 Hierarchical Score Aggregation
Final quality scores compute through bottom-up weighted aggregation respecting rubric hierarchy while tracking reliability via standard deviation propagation. Let $R$ be the rubric root, $C(n)$ denote node $n$’s children, and $w(n)$ denote assigned weight.
Leaf Assessment and Reliability
For leaf nodes $\ell$, multiple assessments $s_{1},s_{2},\ldots,s_{m}$ aggregate as:
$S(\ell)=\bar{s}=\frac{1}{m}\sum_{i=1}^{m}s_{i}$
Reliability quantified via standard deviation:
$\sigma_{\ell}=\sqrt{\frac{1}{m-1}\sum_{i=1}^{m}(s_{i}-\bar{s})^{2}}$
Lower deviations indicate higher judge consensus and assessment reliability.
Hierarchical Propagation
For internal nodes $n$ with children $C(n)=\{c_{1},\ldots,c_{k}\}$ having scores $S(c_{i})$, weights $w(c_{i})$, and deviations $\sigma_{c_{i}}$:
$S(n)=\frac{\sum_{i=1}^{k}w(c_{i})\cdot S(c_{i})}{\sum_{i=1}^{k}w(c_{i})},\quad\sigma_{n}=\frac{\sqrt{\sum_{i=1}^{k}w(c_{i})^{2}\sigma_{c_{i}}^{2}}}{\sum_{i=1}^{k}w(c_{i})}$
This ensures uncertainty propagates appropriately, with weights determining score contribution and reliability influence.
Final Assessment
Final repository score $S(R)\in[0,1]$ with deviation $\sigma_{R}$ represents comprehensive weighted aggregation propagating assessments and reliability upward. Lower $\sigma_{R}$ indicates higher methodology reliability. This dual-metric approach provides quality assessment with confidence bounds, enabling nuanced interpretation. Weights reflect component importance based on criticality and complexity while maintaining statistical rigor in uncertainty quantification.
Repository Language LOC System Score (%) Coverage Improvement (%)
All-Hands-AI–OpenHands Python 229,909 DeepWiki 73.04 ± 2.54 54/67
CodeWiki-kimi-k2 80.20 ± 3.26 60/67 +7.16
CodeWiki-sonnet-4 82.45 ± 2.65 59/67 +9.41
sveltejs–svelte JavaScript 124,576 DeepWiki 68.51 ± 3.31 76/96
CodeWiki-kimi-k2 70.41 ± 3.39 80/96 +1.90
CodeWiki-sonnet-4 71.96 ± 3.73 80/96 +3.45
puppeteer–puppeteer TypeScript 136,302 DeepWiki 64.46 ± 3.72 60/82
CodeWiki-kimi-k2 81.33 ± 3.33 67/82 +16.87
CodeWiki-sonnet-4 83.00 ± 3.37 74/82 +18.54
Unity-Technologies–ml-agents C# 86,106 DeepWiki 74.80 ± 3.69 39/46
CodeWiki-kimi-k2 69.64 ± 5.40 40/46 -5.16
CodeWiki-sonnet-4 79.78 ± 5.02 42/46 +4.98
elastic–logstash Java 117,485 DeepWiki 54.80 ± 4.10 38/57
CodeWiki-kimi-k2 56.89 ± 2.96 37/57 +2.09
CodeWiki-sonnet-4 57.90 ± 3.43 38/57 +3.10
wazuh–wazuh C 1,446,730 DeepWiki 68.68 ± 4.74 39/46
CodeWiki-kimi-k2 58.15 ± 6.20 32/46 -10.53
CodeWiki-sonnet-4 64.17 ± 5.44 34/46 -4.51
electron–electron C++ 184,234 DeepWiki 44.10 ± 3.12 54/92
CodeWiki-kimi-k2 36.99 ± 3.98 43/92 -7.11
CodeWiki-sonnet-4 42.30 ± 3.26 48/92 -1.80
Average DeepWiki 64.06 ± 3.60
CodeWiki-kimi-k2 64.80 ± 4.07 +0.74
CodeWiki-sonnet-4 68.79 ± 3.84 +4.73
Table 1: Documentation generation performance comparison across diverse repositories. Scores represent weighted averages with standard deviations indicating assessment consensus. Coverage shows satisfied leaf criteria out of total requirements. Blue-highlighted rows show CodeWiki-sonnet-4 results, while yellow-highlighted rows show CodeWiki-kimi-k2 results. Improvement column shows relative gain over DeepWiki baseline.
4 Experiment
We conduct comprehensive experiments to evaluate CodeWiki’s effectiveness in generating repository-level documentation across diverse programming languages and project types. Our evaluation is structured around three core research questions that address the fundamental challenges of automated repository documentation generation.
4.1 Research Questions
Our experimental evaluation is guided by the following research questions:
RQ1: Documentation Quality and Coverage How does CodeWiki compare to existing state-of-the-art repository documentation systems in terms of overall documentation quality and requirement coverage?
RQ2: Cross-Language Generalization Does CodeWiki demonstrate consistent performance across diverse programming languages, particularly comparing high-level scripting languages (Python, JavaScript, TypeScript) versus systems programming languages (C, C++, C#, Java)?
RQ3: Scalability and Reliability How does CodeWiki’s hierarchical decomposition approach perform across repositories of varying sizes and complexity, and what is the reliability of our evaluation methodology?
4.2 Baseline
To the best of our knowledge, CodeWiki represents the first open-source framework specifically designed for comprehensive repository-level documentation generation. The absence of comparable open-source solutions necessitates comparison with closed-source alternatives. We select DeepWiki  (DeepWiki, 2025) as our primary baseline, as it shares our goal of automated repository-level documentation generation and has demonstrated effectiveness in industrial applications.
4.3 Experimental Setup
Repository Selection
We carefully curate a comprehensive benchmark consisting of 21 open-source repositories that represent diverse programming languages, project scales, and application domains. The selection criteria prioritize: (1) language diversity to test cross-language generalization, (2) varying codebase sizes from 16K to 1.7M lines of code, (3) active maintenance with comprehensive official documentation for rubric generation, and (4) diverse application domains including web frameworks, automation tools, machine learning platforms, and system utilities.
Table 1 presents the main evaluation set of 7 repositories with detailed analysis, while Table 4 in Appendix D provides results for 14 additional repositories that further validate our approach across diverse codebases.
Model Configuration
We evaluate CodeWiki with Claude Sonnet 4 (CodeWiki-sonnet-4) and Kimi K2 Instruct (CodeWiki-kimi-k2). Rubric generation uses multiple models (Claude Sonnet 4, Gemini 2.5 Pro, Kimi K2 Instruct) to reduce bias. Evaluation employs three judge models (Gemini 2.5 Flash, GPT OSS 120B, Kimi K2 Instruct) for independent assessments (details in Appendix E).
Evaluation Protocol
For each repository, we execute the complete CodeWiki pipeline and generate comprehensive documentation including architectural overviews, component descriptions, and usage guides. The evaluation process follows our hierarchical rubric methodology: (1) generating repository-specific rubrics from official documentation, (2) performing agentic assessment of generated documentation, and (3) computing weighted scores with uncertainty quantification through standard deviation tracking.
4.4 Results and Analysis
4.4.1 RQ1: Documentation Quality and Coverage
Table 1 presents comprehensive evaluation results addressing documentation quality. CodeWiki-sonnet-4 achieves an average quality score of 68.79%, representing a 4.73 percentage point improvement over DeepWiki’s 64.06%. Notably, CodeWiki-kimi-k2 demonstrates competitive performance at 64.80%, closely matching DeepWiki despite using an open-source model. This validates our framework’s effectiveness across different LLM backends.
The coverage metrics reveal that CodeWiki consistently satisfies more leaf-level requirements than the baseline. CodeWiki-sonnet-4 outperforms DeepWiki in 5 out of 7 repositories, with particularly strong improvements in TypeScript (83.00% vs. 64.46%, +18.54%), Python (82.45% vs. 73.04%, +9.41%), and JavaScript (71.96% vs. 68.51%, +3.45%). CodeWiki-kimi-k2 shows notable gains in TypeScript (+16.87%) and Python (+7.16%), demonstrating that our hierarchical approach enables strong performance even with open-source models.
Additional evaluation on 14 diverse repositories in Appendix D further validates these findings, with CodeWiki-kimi-k2 achieving an average improvement of 1.25% and substantial gains on large-scale projects including Trino (+29.53%) and Storybook (+11.42%).
Answer to RQ1: CodeWiki demonstrates superior overall documentation quality and coverage compared to existing systems. With proprietary models, it achieves 4.73% improvement; with open-source models, it matches baseline performance while offering transparency and accessibility benefits.
4.4.2 RQ2: Cross-Language Generalization
Figure 3: Performance comparison by programming language categories. CodeWiki-sonnet-4 shows strong improvements in high-level scripting languages (+10.47%) and managed/enterprise languages (+4.04%), while systems programming languages present challenges for both systems.
Our analysis reveals distinct performance patterns across programming language categories, as visualized in Figure 3. CodeWiki-sonnet-4 demonstrates exceptional performance on high-level scripting languages, achieving 79.14% average across Python, JavaScript, and TypeScript repositories—a 10.47 percentage point improvement over DeepWiki’s 68.67%. CodeWiki-kimi-k2 achieves 77.31% on these languages, showing strong performance with open-source models (+8.64% over DeepWiki).
For managed/enterprise languages (C#, Java), CodeWiki-sonnet-4 achieves competitive performance with an average improvement of 4.04% (68.84% vs. 64.80%), while CodeWiki-kimi-k2 shows 63.27% (-1.53%). Systems programming languages present unique challenges where both frameworks exhibit similar patterns: CodeWiki-sonnet-4 achieves 53.24% compared to DeepWiki’s 56.39%, while CodeWiki-kimi-k2 obtains 47.58%. This suggests that specialized parsing strategies may benefit all approaches for handling complex constructs like pointer operations and template metaprogramming.
Answer to RQ2: CodeWiki demonstrates strong cross-language generalization with exceptional performance on high-level languages (79.14% with proprietary models, 77.31% with open-source models) and consistent improvements on managed languages, while systems programming languages present similar challenges across all frameworks and model types.
4.4.3 RQ3: Scalability and Reliability
Repository size analysis demonstrates CodeWiki’s scalability across 86K-1.4M LOC. For high-level languages, CodeWiki maintains strong performance across scales: Unity ML-Agents (86K LOC) achieves 79.78%, while OpenHands (230K LOC) reaches 82.45%, demonstrating that hierarchical decomposition successfully addresses context limitations. Both CodeWiki and DeepWiki exhibit similar behavioral patterns across repository sizes within language families, with performance variations primarily attributable to language-specific parsing complexity rather than scale. Dynamic delegation automatically adapts to complexity levels, maintaining bounded processing requirements.
Reliability analysis through standard deviations (3.84% CodeWiki vs. 3.60% DeepWiki) validates our multi-model assessment methodology. Lower deviations for high-level languages (Python: 2.65%, JavaScript: 3.73%) versus systems languages (C: 5.44%, C++: 3.26%) confirm evaluation consistency correlates with language characteristics rather than repository scale.
Answer to RQ3: CodeWiki’s hierarchical decomposition enables consistent scalability across repository sizes from 86K to 1.4M LOC, with performance variations primarily attributable to language-specific characteristics rather than scale limitations. Both frameworks exhibit similar patterns across different repository sizes within language families.
5 Conclusion
We presented CodeWiki, an open-source semi-agentic framework addressing automated repository-level documentation through hierarchical decomposition and recursive processing. CodeWiki achieves 68.79% quality with proprietary models (Claude Sonnet 4) and 64.80% with open-source alternatives (Kimi K2 Instruct), both outperforming or matching DeepWiki’s 64.06%, demonstrating strong performance regardless of underlying model choice.
Results show performance variations correlate primarily with language characteristics rather than repository size or model type. High-level languages show consistent improvements across scales and models, while systems languages face challenges regardless of configuration.
Future directions include: developing specialized parsing modules for improved systems language support, extending to multi-version documentation tracking codebase evolution, and leveraging comprehensive documentation for downstream tasks like code migration. By releasing CodeWiki as open source with support for both proprietary and open-source LLMs, we accelerate research and enable the community to build upon our hierarchical processing approach with accessible technology choices.
Limitations
While CodeWiki demonstrates strong performance across multiple languages and repository scales, two main limitations warrant discussion. First, our framework exhibits reduced effectiveness on systems programming languages (C, C++) compared to high-level languages, achieving 53.24% versus 79.14% average scores. This gap stems from the inherent complexity of low-level constructs such as pointer operations, manual memory management, and template metaprogramming, which challenge both AST parsing and semantic understanding.
Second, our evaluation methodology, while more comprehensive than traditional metrics, relies on LLM-based assessment which may introduce model-specific biases despite our multi-model consensus approach. The average standard deviation of 3.84% indicates reasonable consistency, yet edge cases may benefit from human expert validation.
Appendices
Appendix A Example Generated Documentation
Figure 4: Example repository-level documentation generated by CodeWiki for the All-Hands-AI–OpenHands repository. The documentation includes a comprehensive overview, architectural diagrams showing the modular event-driven architecture, and hierarchical navigation of core components including the agent system, implementations, and event processing modules.
Figure 4 shows an example of repository-level documentation generated by CodeWiki for the All-Hands-AI–OpenHands repository. The documentation provides a comprehensive overview including the repository’s purpose, end-to-end architecture with visual diagrams, and hierarchical navigation of core components. This demonstrates CodeWiki’s ability to synthesize both textual explanations and visual artifacts that capture architectural patterns and module relationships, enabling developers to quickly understand the system’s structure and design decisions.
Appendix B Background and Related Work
B.1 Code Summarization and Documentation
B.1.1 Existing Approaches
Advances in natural language processing (NLP) have enabled significant progress in the task of code summarization. Existing methods  (Khan and Uddin, 2023; Poudel et al., 2024; Makharev and Ivanov, 2025; Lomshakov et al., 2024; Ahmed and Devanbu, 2023; Choi et al., 2023; Luo et al., 2024; Yang et al., 2025) primarily focus on function-level documentation, and can broadly be divided into two categories based on the type of context they exploit (see Table 2 for a comprehensive comparison):
Local-Level Context
Early neural approaches established the foundation for automated documentation. CodeBERT (Feng et al., 2020) demonstrated effective bimodal training on code and natural language, while more recent LLM applications have shown substantial improvements. For instance, (Khan and Uddin, 2023) reported 20.6 BLEU scores across six languages using GPT-3, an 11.2% improvement over prior methods. DocuMint (Poudel et al., 2024) further explored the use of small language models for Python docstring generation with promising results. However, these approaches primarily rely on local code context and do not leverage repository-level information to improve documentation quality.
Codebase-Level Context
Recognizing the limitations of function-only context, recent work has explored broader codebase information to enrich documentation.  (Makharev and Ivanov, 2025) demonstrated that incorporating class- and repository-level context can substantially improve summary quality through few-shot learning and retrieval-augmented generation. ProConSuL (Lomshakov et al., 2024) incorporated project-level information via compiler analysis and call graphs, employing a two-phase training procedure to reduce hallucinations. RepoAgent (Luo et al., 2024) introduced comprehensive LLM-powered repository documentation, though it was limited in content diversity and architectural depth. DocAgent (Yang et al., 2025) advanced the field through multi-agent collaboration with topological processing, employing specialized agents that respect code dependencies to improve completeness and accuracy.
$\hookrightarrow$ Limitations and Our Motivation: Despite this progress, most approaches still target isolated snippets (e.g., functions or classes) and thus fail to capture the underlying architecture of the repository, as highlighted in Table 2. For a developer onboarding a completely new codebase, the primary challenge is not understanding what each individual function or class does, but rather grasping the high-level overview of the system, its dataflow and how different components interact. This underscores the need for an automatic methodology capable of generating documentation that spans from low-level details to high-level repository-wide perspectives, covering critical features and providing visual representations that support developer comprehension. Our approach, CodeWiki, addresses these limitations by generating true repository-level documentation while leveraging full codebase context and supporting multiple programming languages.
B.1.2 Existing Evaluation Methods
Most existing works evaluate the quality of generated documentation using textual similarity metrics such as BLEU or ROUGE. These metrics measure the overlap between generated text and reference ground-truth documentation in terms of lexical or token-level similarity.
$\hookrightarrow$ Limitations and Our Motivation: While these methods provide a lightweight quantitative measure, they often fail to capture deeper semantic aspects of the documentation, such as clarity, correctness, and completeness in explaining the underlying code behavior. Recent studies (Roy et al., 2021; Evtikhiev et al., 2023) have shown that traditional metrics like BLEU and ROUGE do not consistently correlate with human judgment of documentation quality. This gap highlights the need for an assessment system that can fairly and accurately evaluate documentation, ensuring that the generated content truly supports developers in comprehending the system.
Table 2: Comparison of different code summarization approaches. Repo-Level Doc. indicates whether the approach generates repository-level documentation output. Codebase Context specifies whether the codebase information is used as input to generate the documentation.
Approach Repo-Level Doc. Codebase Context Multilingual DocuMint ASAP ProConSul DocAgent RepoAgent CodeWiki
B.2 Rubric-based evaluation
A repository can be described through multiple representations, each using different structures and presentation styles while sharing the same underlying architecture. This diversity makes it difficult to evaluate open-ended outputs automatically, as no single metric can reliably evaluate which version is better. Although human evaluation provides trustworthy results, it is resource-intensive and slow. To overcome these limitations and facilitate rapid assessment, many benchmarks (Starace et al., 2025; Lin et al., 2025; Sirdeshmukh et al., 2025; Arora et al., 2025) often employ large language models (LLMs) as automated evaluators, using task-specific rubrics to judge the quality of generated outputs. In our CodeWikiBench benchmark, we similarly construct a set of hierarchical rubrics based on the official documentation, providing a structured and fine-grained framework to evaluate the quality and completeness of generated repository-level documentation.
Appendix C Rubric Reliability Assessment
To ensure the quality and trustworthiness of automatically generated evaluation rubrics, we developed a comprehensive reliability assessment framework that quantifies the consistency and structural integrity of rubrics produced by different language models. This assessment methodology provides empirical measures of rubric reliability across multiple dimensions.
Given a set of $n$ rubrics $\{R_{1},R_{2},\ldots,R_{n}\}$ produced by different models for the same documentation corpus, we compute pairwise consistency scores across both semantic and structural dimensions.
C.1 Semantic Consistency Analysis
We evaluate semantic consistency by extracting all requirement texts from each rubric and compute semantic similarity using distributed representations. Let $T_{i}=\{t_{1}^{(i)},t_{2}^{(i)},\ldots,t_{m_{i}}^{(i)}\}$ denote the set of requirement texts extracted from rubric $R_{i}$, where $m_{i}$ represents the total number of requirements in $R_{i}$.
For each requirement text $t_{k}^{(i)}\in T_{i}$, we identify its best semantic match in $T_{j}$ using cosine similarity between embeddings:
$\text{sim}_{\text{best}}(t_{k}^{(i)},T_{j})=\max_{t_{\ell}^{(j)}\in T_{j}}\cos(\mathbf{e}(t_{k}^{(i)}),\mathbf{e}(t_{\ell}^{(j)}))$ (1)
where $\mathbf{e}(\cdot)$ represents the embedding function and $\cos(\cdot,\cdot)$ denotes cosine similarity. The bidirectional semantic similarity between rubrics is then:
$\text{Sim}_{\text{semantic}}(R_{i},R_{j})$ $=\frac{1}{m_{i}+m_{j}}\Bigg(\sum_{k=1}^{m_{i}}\text{sim}_{\text{best}}(t_{k}^{(i)},T_{j})$
$\quad+\sum_{\ell=1}^{m_{j}}\text{sim}_{\text{best}}(t_{\ell}^{(j)},T_{i})\Bigg)$ (2)
This formulation ensures that both the coverage of requirements from $R_{i}$ in $R_{j}$ and vice versa are considered, providing a balanced measure of semantic overlap.
C.2 Structural Consistency Analysis
Structural consistency evaluates the architectural similarity between rubrics, focusing on hierarchical organization rather than content semantics. We define a comprehensive set of structural features for each rubric:
  • Maximum depth ($d$): The deepest level in the rubric hierarchy
  • Total items ($N$): The total number of rubric items across all levels
  • Weight distribution ($W$): The frequency distribution of importance weights
The structural similarity between rubrics $R_{i}$ and $R_{j}$ combines multiple normalized metrics:
$\text{Sim}_{\text{depth}}(R_{i},R_{j})=1-\frac{|d_{i}-d_{j}|}{\max(d_{i},d_{j},1)}$ (3)
$\text{Sim}_{\text{items}}(R_{i},R_{j})=1-\frac{|N_{i}-N_{j}|}{\max(N_{i},N_{j},1)}$ (4)
For weight distribution similarity, we compute the overlap between normalized probability distributions:
$\text{Sim}_{\text{weights}}(R_{i},R_{j})=\sum_{w\in W_{i}\cup W_{j}}\min(P_{i}(w),P_{j}(w))$ (5)
where $P_{i}(w)$ represents the normalized frequency of weight $w$ in rubric $R_{i}$.
The overall structural similarity is computed as:
$\text{Sim}_{\text{structural}}(R_{i},R_{j})$ $=\frac{1}{3}\Big(\text{Sim}_{\text{depth}}(R_{i},R_{j})$
$\quad+\text{Sim}_{\text{items}}(R_{i},R_{j})$ (6)
$\quad+\text{Sim}_{\text{weights}}(R_{i},R_{j})\Big)$ (7)
C.3 Overall Reliability Score
The overall reliability score aggregates consistency measures across all model pairs. For $n$ models, we compute $\binom{n}{2}$ pairwise comparisons and derive summary statistics:
$\text{Score}^{semantic}_{\text{reliability}}=\frac{1}{\binom{n}{2}}\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\text{Sim}_{\text{semantic}}(R_{i},R_{j})$ (8)
$\text{Score}^{structural}_{\text{reliability}}=\frac{1}{\binom{n}{2}}\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\text{Sim}_{\text{structural}}(R_{i},R_{j})$ (9)
Table 3 presents the reliability assessment results across our benchmark repositories, demonstrating consistent rubric generation with average semantic reliability of 73.65% and structural reliability of 70.84%.
Table 3: Rubric Reliability Assessment: Semantic and Structural Consistency Scores
Repository Semantic (%) Structural (%) All-Hands-AI–OpenHands 73.64 68.82 sveltejs–svelte 75.56 68.88 puppeteer–puppeteer 72.97 76.36 Unity-Technologies–ml-agents 75.95 71.22 elastic–logstash 73.13 68.22 wazuh–wazuh 72.20 67.62 electron–electron 72.12 74.79 Average 73.65 70.84
Appendix D Complementary Experimental Results
To further validate CodeWiki’s effectiveness across diverse programming languages and repository types, we conducted additional experiments on 14 repositories spanning 7 programming languages. Table 4 presents comprehensive results using CodeWiki-kimi-k2 (documentation generated using Kimi K2 Instruct) compared against the DeepWiki baseline.
Repository Language LOC System Score (%) Coverage Improvement (%)
chartjs–Chart.js JavaScript 90,925 DeepWiki 88.71 ± 3.29 57/61
CodeWiki-kimi-k2 84.01 ± 3.55 54/61 -4.70
marktext–marktext JavaScript 61,574 DeepWiki 81.64 ± 4.67 50/57
CodeWiki-kimi-k2 82.49 ± 2.51 51/57 +0.85
storybookjs–storybook TypeScript 364,606 DeepWiki 72.07 ± 3.59 64/78
CodeWiki-kimi-k2 83.49 ± 2.31 69/78 +11.42
mermaid-js–mermaid TypeScript 157,160 DeepWiki 73.81 ± 4.05 59/67
CodeWiki-kimi-k2 71.52 ± 5.36 57/67 -2.29
qmk–qmk_firmware C 1,099,824 DeepWiki 81.23 ± 2.62 94/110
CodeWiki-kimi-k2 72.01 ± 2.63 88/110 -9.22
sumatrapdfreader–sumatrapdf C 1,699,044 DeepWiki 55.81 ± 6.29 31/40
CodeWiki-kimi-k2 58.92 ± 5.18 31/40 +3.11
x64dbg–x64dbg C++ 207,671 DeepWiki 64.42 ± 2.07 62/107
CodeWiki-kimi-k2 58.01 ± 3.58 56/107 -6.41
nlohmann–json C++ 119,542 DeepWiki 66.06 ± 3.08 33/57
CodeWiki-kimi-k2 61.28 ± 2.35 30/57 -4.78
FluentValidation–FluentValidation C# 16,452 DeepWiki 76.48 ± 2.33 41/48
CodeWiki-kimi-k2 70.03 ± 2.90 38/48 -6.45
git-ecosystem–git-credential-manager C# 40,812 DeepWiki 87.03 ± 2.58 66/70
CodeWiki-kimi-k2 82.78 ± 0.90 61/70 -4.25
material-components–material-components-android Java 244,092 DeepWiki 70.99 ± 2.92 79/96
CodeWiki-kimi-k2 79.16 ± 4.32 89/96 +8.17
trinodb–trino Java 1,535,202 DeepWiki 55.27 ± 3.76 48/68
CodeWiki-kimi-k2 84.80 ± 2.84 62/68 +29.53
RasaHQ–rasa Python 147,711 DeepWiki 80.94 ± 3.15 64/70
CodeWiki-kimi-k2 81.10 ± 2.69 64/70 +0.16
microsoft–graphrag Python 30,390 DeepWiki 76.34 ± 3.28 48/56
CodeWiki-kimi-k2 79.29 ± 2.96 50/56 +2.95
Average DeepWiki 73.62 ± 3.41
CodeWiki-kimi-k2 74.87 ± 3.15 +1.25
Table 4: Complementary documentation generation performance evaluation across 14 additional repositories. Results demonstrate CodeWiki’s consistent effectiveness across diverse codebases, with notable improvements in Java repositories (trinodb–trino: +29.53%, material-components-android: +8.17%) and TypeScript (storybookjs–storybook: +11.42%).
These complementary results reinforce our findings from the main experiments, showing an average improvement of 1.25% across 14 diverse repositories. Notably, CodeWiki achieves substantial gains in several large-scale projects, including a remarkable 29.53% improvement on the Trino database engine (1.5M LOC) and 11.42% on the Storybook UI development platform (365K LOC). The results further validate that CodeWiki’s hierarchical approach generalizes well across different repository characteristics and programming paradigms.
Appendix E Implementation Details
E.1 Model Configurations
Documentation Generation Agent
We evaluate CodeWiki with two model configurations:
  • CodeWiki-sonnet-4: Claude Sonnet 4 (claude-sonnet-4-20250514)
    • Temperature: 0.0
    • Max tokens: 32768
    • Context window: 200K tokens
  • CodeWiki-kimi-k2: Kimi K2 Instruct (kimi-k2-instruct-0905)
    • Temperature: 0.0
    • Max tokens: 32768
    • Context window: 200K tokens
Rubric Generation Agents
  • Model 1: Claude Sonnet 4 (Temperature: 0.0)
  • Model 2: Gemini 2.5 Pro (Temperature: 0.0)
  • Model 3: Kimi K2 Instruct (Temperature: 0.0)
  • Max tokens per generation: 32768
Judge Agents
  • Judge 1: Gemini 2.5 Flash (Temperature: 0.0)
  • Judge 2: GPT OSS 120B (Temperature: 0.0)
  • Judge 3: Kimi K2 Instruct (Temperature: 0.0)
  • Binary scoring: 0 (not satisfied) or 1 (satisfied)
E.2 Processing Parameters
  • Module decomposition threshold: 32768 tokens per leaf module
  • Maximum delegation depth: 3 levels

### v2

@@
-CodeWiki: Automated Repository-Level Documentation at Scale
+CodeWiki: Evaluating AI’s Ability to Generate Holistic Documentation for Large-Scale Codebases
 Abstract
-Developers spend nearly 58% of their time understanding codebases, yet maintaining comprehensive documentation remains challenging due to complexity and manual effort.
-While recent Large Language Models (LLMs) show promise for function-level documentation, they fail at the repository level, where capturing architectural patterns and cross-module interactions is essential.
-We introduce CodeWiki, the first open-source framework for holistic repository-level documentation across seven programming languages.
-CodeWiki employs three innovations: (i) hierarchical decomposition that preserves architectural context, (ii) recursive agentic processing with dynamic delegation, and (iii) synthesis of textual and visual artifacts including architecture diagrams and data flows.
-We also present CodeWikiBench, the first repository-level documentation benchmark with multi-level rubrics and agentic assessment.
-CodeWiki achieves 68.79% quality score with proprietary models and 64.80% with open-source alternatives, outperforming existing closed-source systems and demonstrating scalable, accurate documentation for real-world repositories.
+Given a large and evolving codebase, the ability to automatically generate a holistic, architecture-aware documentation that captures not only individual functions but also their cross-file, cross-module, and system-level interactions remains an open challenge.
+Comprehensive documentation is essential for long-term software maintenance and team collaboration, yet current automated approaches still fail to model the rich semantic dependencies and architectural structures that define real-world software systems.
+We present CodeWiki, a unified framework for automated repository-level documentation across seven programming languages.
+CodeWiki introduces three key innovations: (i) hierarchical decomposition that preserves architectural context across multiple levels of granularity, (ii) recursive multi-agent processing with dynamic task delegation for scalable generation, and (iii) multi-modal synthesis that integrates textual descriptions with visual artifacts such as architecture diagrams and data-flow representations.
+To enable rigorous evaluation, we introduce CodeWikiBench, a comprehensive benchmark featuring multi-dimensional rubrics and LLM-based assessment protocols.
+Experimental results demonstrate that CodeWiki achieves a 68.79% quality score with proprietary models, outperforming the closed-source DeepWiki baseline (64.06%) by 4.73%, with particularly strong improvements on high-level scripting languages (+10.47%).
+We open-source CodeWiki to foster future research and community adoption.
+[…]
+[…]
 1 Introduction
@@
 In the rapidly evolving landscape of software development, maintaining comprehensive and up-to-date documentation has become increasingly challenging as codebases grow in size and complexity.
-Studies indicate that developers spend approximately 58% of their working time on program comprehension activities (Xia et al., 2018), highlighting the critical importance of accessible and accurate documentation.
-Despite documentation being recognized as essential for software maintenance and collaboration (de Souza et al., 2005; Zhi et al., 2015; Chen and Huang, 2009), the manual creation and maintenance of repository-level documentation remain a labor-intensive and often neglected aspect of software engineering practice (McBurney et al., 2018).
-The emergence of Large Language Models (LLMs)  (Zhang et al., 2023; He et al., 2025b) has opened opportunities for automating documentation generation tasks.
+Recent industry surveys reveal that documentation represents a critical bottleneck in modern development workflows: approximately 31% of developers currently rely heavily on AI for documenting code, while 25% use AI extensively for creating and maintaining documentation—making documentation tasks among the highest priorities for AI integration over the next 3-5 years [25].
+Notably, while developers express high satisfaction with AI tools for searching and learning, they indicate plans to substantially increase AI adoption for documentation and testing tasks, suggesting current tools remain insufficient for these critical needs.
+This widespread focus on AI-driven documentation reflects its dual burden: while essential for software maintenance and collaboration [6, 31, 3], manual creation and maintenance remain labor-intensive and often neglected [19].
+The emergence of Large Language Models (LLMs) [30, 12] has opened opportunities for automating documentation generation.
 However, current approaches still face two fundamental limitations.
 Limitations of Scalability at the Repository Level
-Recent advances in code understanding and natural language generation have demonstrated promising results in function-level and file-level documentation tasks (Feng et al., 2020; Khan and Uddin, 2023; Poudel et al., 2024; Makharev and Ivanov, 2025; Lomshakov et al., 2024; Ahmed and Devanbu, 2023; Choi et al., 2023; Luo et al., 2024; Yang et al., 2025) However, existing approaches face significant challenges when scaling to repository-level documentation.
-Unlike function-level documentation that focuses on individual components, repository-level documentation task is to capture architectural patterns, cross-module interactions, data flows, and system-wide design decisions (Rai et al., 2022; Treude et al., 2020).
-Furthermore, the hierarchical nature of software systems requires documentation that can serve multiple audiences, from high-level stakeholders seeking architectural overviews to developers requiring detailed implementation guidance (Nassif and Robillard, 2025).
+Recent advances in code understanding and natural language generation have demonstrated promising results in function-level and file-level documentation tasks [9, 13, 21, 18, 16, 1, 4, 17, 29] However, existing approaches face significant challenges when scaling to repository-level documentation.
+Unlike function-level documentation that focuses on individual components, repository-level documentation task is to capture architectural patterns, cross-module interactions, data flows, and system-wide design decisions [22, 27].
+Furthermore, the hierarchical nature of software systems requires documentation that can serve multiple audiences, from high-level stakeholders seeking architectural overviews to developers requiring detailed implementation guidance [20].
 $\hookrightarrow$ To address these challenges, we present CodeWiki, the first open-source framework specifically designed to generate holistic, structured repository-level documentation that includes diverse content types such as system architecture diagrams, data flow visualizations, and sequence diagrams while scaling effectively for projects of arbitrary size.
@@
 Despite progress, systematic benchmarks remain lacking.
-Function-level and file-level work relies on general-purpose metrics (BLEU, ROUGE) that fail to capture documentation quality nuances (Roy et al., 2021; Evtikhiev et al., 2023).
+Function-level and file-level work relies on general-purpose metrics (BLEU, ROUGE) that fail to capture documentation quality nuances [23, 8].
 Repository-level evaluation is particularly challenging as documentation admits multiple valid structures sharing underlying architectural content, highlighting the need for robust assessment beyond surface-level similarity.
@@
 Dependency Graph Construction
-Following DocAgent (Yang et al., 2025), we employ Tree-Sitter parsers to extract ASTs, systematically identifying functions, methods, classes, structs, modules, and interdependencies: function calls, class inheritance, attribute access, and module imports.
-We normalize these to a unified depends_on relation for cross-language generalization, creating directed graph $G=(V,E)$ where edge $(u,v)\in E$ indicates component $u$ depends on $v$ (Liu et al., 2023).
+Following DocAgent [29], we employ Tree-Sitter parsers to extract ASTs, systematically identifying functions, methods, classes, structs, modules, and interdependencies: function calls, class inheritance, attribute access, and module imports.
+We normalize these to a unified depends_on relation for cross-language generalization, creating directed graph $G=(V,E)$ where edge $(u,v)\in E$ indicates component $u$ depends on $v$ [15].
 Entry Point Identification and Decomposition
@@
 Output: Complete repository documentation
-Initialize documentation workspace;
-for each leaf module $m$ in $T$ do
-   agent $\leftarrow$ CreateAgent($m$, $G$, $T$);
- while $m$ not fully processed do
-      agent.GenerateDoc($m$.name);
-    if delegation requested then
-         sub $\leftarrow$ agent.GetDelegationSpecs();
-         UpdateModuleTree($T$, $m$, sub);
-       foreach new leaf $m^{\prime}$ in sub do
-            Recursively process $m^{\prime}$;
-        ReviseDoc($m$.name);
-ProcessParentModules($T$);
-return Repository documentation;
+1 Initialize documentation workspace;
+2 for each leaf module $m$ in $T$ do
+3   agent $\leftarrow$ CreateAgent($m$, $G$, $T$);
+4 while $m$ not fully processed do
+5      agent.GenerateDoc($m$.name);
+6    if delegation requested then
+7         sub $\leftarrow$ agent.GetDelegationSpecs();
+8         UpdateModuleTree($T$, $m$, sub);
+9       foreach new leaf $m^{\prime}$ in sub do
+10            Recursively process $m^{\prime}$;
+11
+12         end foreach
+13        ReviseDoc($m$.name);
+14
+15      end if
+16
+17   end while
+18
+19 end for
+20ProcessParentModules($T$);
+21 return Repository documentation;
 Algorithm 1 Recursive Module Documentation
@@
 3 The First Benchmarking System - CodeWikiBench
+Repository-level documentation evaluation presents unique challenges due to complexity and hierarchical organization.
+Unlike code generation assessable through automated metrics, documentation quality requires nuanced evaluation of completeness, accuracy, and coherence across abstraction levels.
+Recent research highlights traditional metrics’ limitations for code tasks [8, 10], while LLM-as-a-Judge paradigms offer new possibilities [28, 5, 11].
+We develop a comprehensive framework synthesizing repository-specific rubrics and employing agentic assessment.
+3.1 Hierarchical Rubric Generation
 Figure 2: Example hierarchical evaluation rubric for the RAGFlow repository.
@@
 The hierarchical organization ensures comprehensive coverage from high-level architectural components down to specific implementation details, with weights reflecting the relative importance of each component for understanding the repository
-Repository-level documentation evaluation presents unique challenges due to complexity and hierarchical organization.
-Unlike code generation assessable through automated metrics, documentation quality requires nuanced evaluation of completeness, accuracy, and coherence across abstraction levels.
-Recent research highlights traditional metrics’ limitations for code tasks (Evtikhiev et al., 2023; Haldar and Hockenmaier, 2024), while LLM-as-a-Judge paradigms offer new possibilities (Wang et al., 2025; Crupi et al., 2025; He et al., 2025a).
-We develop a comprehensive framework synthesizing repository-specific rubrics and employing agentic assessment.
-3.1 Hierarchical Rubric Generation
-Following (Starace et al., 2025), we synthesize evaluation rubrics tailored for each repository, illustrated in Figure 2.
+Following [26], we synthesize evaluation rubrics tailored for each repository, illustrated in Figure 2.
 We collect official documentation from selected open-source projects and hierarchically parse into structured JSON format, leveraging directory structure and markdown syntax.
@@
 The absence of comparable open-source solutions necessitates comparison with closed-source alternatives.
-We select DeepWiki  (DeepWiki, 2025) as our primary baseline, as it shares our goal of automated repository-level documentation generation and has demonstrated effectiveness in industrial applications.
+We select DeepWiki [7] as our primary baseline, as it shares our goal of automated repository-level documentation generation and has demonstrated effectiveness in industrial applications.
 4.3 Experimental Setup
-Repository Selection
+4.3.1 Repository Selection
 We carefully curate a comprehensive benchmark consisting of 21 open-source repositories that represent diverse programming languages, project scales, and application domains.
-The selection criteria prioritize: (1) language diversity to test cross-language generalization, (2) varying codebase sizes from 16K to 1.7M lines of code, (3) active maintenance with comprehensive official documentation for rubric generation, and (4) diverse application domains including web frameworks, automation tools, machine learning platforms, and system utilities.
+The selection criteria prioritize:
+  1. 1.
+    Language diversity to test cross-language generalization
+  2. 2.
+    Varying codebase sizes from 16K to 1.7M lines of code
+  3. 3.
+    Active maintenance with comprehensive official documentation for rubric generation
+  4. 4.
+    Diverse application domains including web frameworks, automation tools, machine learning platforms, and system utilities
 Table 1 presents the main evaluation set of 7 repositories with detailed analysis, while Table 4 in Appendix D provides results for 14 additional repositories that further validate our approach across diverse codebases.
-Model Configuration
-We evaluate CodeWiki with Claude Sonnet 4 (CodeWiki-sonnet-4) and Kimi K2 Instruct (CodeWiki-kimi-k2).
-Rubric generation uses multiple models (Claude Sonnet 4, Gemini 2.5 Pro, Kimi K2 Instruct) to reduce bias.
-Evaluation employs three judge models (Gemini 2.5 Flash, GPT OSS 120B, Kimi K2 Instruct) for independent assessments (details in Appendix E).
-Evaluation Protocol
+4.3.2 Model Configuration
+We evaluate CodeWiki with two different LLM backends:
+  • CodeWiki-sonnet-4: Uses Claude Sonnet 4 (proprietary model)
+  • CodeWiki-kimi-k2: Uses Kimi K2 Instruct (open-source model)
+For rubric generation, we employ multiple models (Claude Sonnet 4, Gemini 2.5 Pro, Kimi K2 Instruct) to reduce single-model bias.
+The evaluation process uses three independent judge models (Gemini 2.5 Flash, GPT OSS 120B, Kimi K2 Instruct) for robust assessment.
+Additional implementation details are provided in Appendix E.
+4.3.3 Evaluation Protocol
 For each repository, we execute the complete CodeWiki pipeline and generate comprehensive documentation including architectural overviews, component descriptions, and usage guides.
-The evaluation process follows our hierarchical rubric methodology: (1) generating repository-specific rubrics from official documentation, (2) performing agentic assessment of generated documentation, and (3) computing weighted scores with uncertainty quantification through standard deviation tracking.
+The evaluation process follows our hierarchical rubric methodology:
+  1. 1.
+    Rubric Generation: Repository-specific rubrics derived from official documentation
+  2. 2.
+    Agentic Assessment: Multiple judge agents evaluate generated documentation against rubric criteria
+  3. 3.
+    Score Aggregation: Weighted scores computed with uncertainty quantification through standard deviation tracking
 4.4 Results and Analysis
@@
 4.4.2 RQ2: Cross-Language Generalization
+Our analysis reveals distinct performance patterns across programming language categories, as visualized in Figure 3.
 Figure 3: Performance comparison by programming language categories.
 CodeWiki-sonnet-4 shows strong improvements in high-level scripting languages (+10.47%) and managed/enterprise languages (+4.04%), while systems programming languages present challenges for both systems.
-Our analysis reveals distinct performance patterns across programming language categories, as visualized in Figure 3.
 CodeWiki-sonnet-4 demonstrates exceptional performance on high-level scripting languages, achieving 79.14% average across Python, JavaScript, and TypeScript repositories—a 10.47 percentage point improvement over DeepWiki’s 68.67%.
@@
 Advances in natural language processing (NLP) have enabled significant progress in the task of code summarization.
-Existing methods  (Khan and Uddin, 2023; Poudel et al., 2024; Makharev and Ivanov, 2025; Lomshakov et al., 2024; Ahmed and Devanbu, 2023; Choi et al., 2023; Luo et al., 2024; Yang et al., 2025) primarily focus on function-level documentation, and can broadly be divided into two categories based on the type of context they exploit (see Table 2 for a comprehensive comparison):
+Existing methods [13, 21, 18, 16, 1, 4, 17, 29] primarily focus on function-level documentation, and can broadly be divided into two categories based on the type of context they exploit (see Table 2 for a comprehensive comparison):
 Local-Level Context
 Early neural approaches established the foundation for automated documentation.
-CodeBERT (Feng et al., 2020) demonstrated effective bimodal training on code and natural language, while more recent LLM applications have shown substantial improvements.
-For instance, (Khan and Uddin, 2023) reported 20.6 BLEU scores across six languages using GPT-3, an 11.2% improvement over prior methods.
-DocuMint (Poudel et al., 2024) further explored the use of small language models for Python docstring generation with promising results.
+CodeBERT [9] demonstrated effective bimodal training on code and natural language, while more recent LLM applications have shown substantial improvements.
+For instance, [13] reported 20.6 BLEU scores across six languages using GPT-3, an 11.2% improvement over prior methods.
+DocuMint [21] further explored the use of small language models for Python docstring generation with promising results.
 However, these approaches primarily rely on local code context and do not leverage repository-level information to improve documentation quality.
@@
 Recognizing the limitations of function-only context, recent work has explored broader codebase information to enrich documentation.
-(Makharev and Ivanov, 2025) demonstrated that incorporating class- and repository-level context can substantially improve summary quality through few-shot learning and retrieval-augmented generation.
-ProConSuL (Lomshakov et al., 2024) incorporated project-level information via compiler analysis and call graphs, employing a two-phase training procedure to reduce hallucinations.
-RepoAgent (Luo et al., 2024) introduced comprehensive LLM-powered repository documentation, though it was limited in content diversity and architectural depth.
-DocAgent (Yang et al., 2025) advanced the field through multi-agent collaboration with topological processing, employing specialized agents that respect code dependencies to improve completeness and accuracy.
+[18] demonstrated that incorporating class- and repository-level context can substantially improve summary quality through few-shot learning and retrieval-augmented generation.
+ProConSuL [16] incorporated project-level information via compiler analysis and call graphs, employing a two-phase training procedure to reduce hallucinations.
+RepoAgent [17] introduced comprehensive LLM-powered repository documentation, though it was limited in content diversity and architectural depth.
+DocAgent [29] advanced the field through multi-agent collaboration with topological processing, employing specialized agents that respect code dependencies to improve completeness and accuracy.
 $\hookrightarrow$ Limitations and Our Motivation: Despite this progress, most approaches still target isolated snippets (e.g., functions or classes) and thus fail to capture the underlying architecture of the repository, as highlighted in Table 2.
@@
 $\hookrightarrow$ Limitations and Our Motivation: While these methods provide a lightweight quantitative measure, they often fail to capture deeper semantic aspects of the documentation, such as clarity, correctness, and completeness in explaining the underlying code behavior.
-Recent studies (Roy et al., 2021; Evtikhiev et al., 2023) have shown that traditional metrics like BLEU and ROUGE do not consistently correlate with human judgment of documentation quality.
+Recent studies [23, 8] have shown that traditional metrics like BLEU and ROUGE do not consistently correlate with human judgment of documentation quality.
 This gap highlights the need for an assessment system that can fairly and accurately evaluate documentation, ensuring that the generated content truly supports developers in comprehending the system.
@@
 Although human evaluation provides trustworthy results, it is resource-intensive and slow.
-To overcome these limitations and facilitate rapid assessment, many benchmarks (Starace et al., 2025; Lin et al., 2025; Sirdeshmukh et al., 2025; Arora et al., 2025) often employ large language models (LLMs) as automated evaluators, using task-specific rubrics to judge the quality of generated outputs.
+To overcome these limitations and facilitate rapid assessment, many benchmarks [26, 14, 24, 2] often employ large language models (LLMs) as automated evaluators, using task-specific rubrics to judge the quality of generated outputs.
 In our CodeWikiBench benchmark, we similarly construct a set of hierarchical rubrics based on the official documentation, providing a structured and fine-grained framework to evaluate the quality and completeness of generated repository-level documentation.
@@
 Table 3: Rubric Reliability Assessment: Semantic and Structural Consistency Scores
-Repository Semantic (%) Structural (%) All-Hands-AI–OpenHands 73.64 68.82 sveltejs–svelte 75.56 68.88 puppeteer–puppeteer 72.97 76.36 Unity-Technologies–ml-agents 75.95 71.22 elastic–logstash 73.13 68.22 wazuh–wazuh 72.20 67.62 electron–electron 72.12 74.79 Average 73.65 70.84
+Repository Semantic (%) Structural (%)
+All-Hands-AI–OpenHands 73.64 68.82
+sveltejs–svelte 75.56 68.88
+puppeteer–puppeteer 72.97 76.36
+Unity-Technologies–ml-agents 75.95 71.22
+elastic–logstash 73.13 68.22
+wazuh–wazuh 72.20 67.62
+electron–electron 72.12 74.79
+Average 73.65 70.84
 Appendix D Complementary Experimental Results

### v3

@@
     Competitive performance on real-world repositories, improving over DeepWiki with both proprietary (Claude Sonnet 4) and open-source (Kimi K2 Instruct) models.
+2 Related Work
+2.1 Code Intelligence and LLMs
+The rapid advancement of Large Language Models (LLMs) has revolutionized code intelligence tasks.
+Early work by von der Mosel et al.
+[47] demonstrated the effectiveness of Transformer-based models for software engineering, leading to specialized code models like CodeBERT [12], GraphCodeBERT [13], CodeT5 [49], CodeT5+ [50], and CodeTF [3].
+These established early transformer-based approaches for code understanding and generation.
+Comprehensive surveys by Niu et al.
+[34] and Hou et al.
+[19] established the importance of domain-specific pre-training, showing consistent improvements over general-purpose language models across code search, summarization, and generation tasks.
+However, recent studies highlight important limitations: Li et al.
+[22] found that models often make confident decisions without sufficient reasoning, while Zheng et al.
+[57] noted that LLM performance remains unstable and requires careful prompt engineering [42].
+For code generation, HumanEval [6] and MBPP [2] established foundational benchmarks, later extended to multiple languages by MultiPL-E [4] and CodeScope [52].
+More recently, CodeMMLU [32] has emerged as a comprehensive multi-task benchmark for assessing code understanding and reasoning capabilities of CodeLLMs.
+2.2 Automated Documentation Generation
+Documentation generation has emerged as a critical LLM application in software engineering.
+Traditional approaches focused on function-level summarization using datasets like CodeSearchNet [15] and FunCom [21].
+While Sun et al.
+[45] found that ChatGPT can generate coherent comments, their quality often falls short of specialized models when evaluated using BLEU and ROUGE metrics.
+Recent research has shifted toward repository-level documentation generation, which requires understanding complex dependencies and architectural context.
+Luo et al.
+[27] introduced RepoAgent, an LLM-powered framework that leverages static analysis and incremental processing.
+DocAgent [53] proposed a multi-agent system with specialized roles (Reader, Searcher, Writer, Verifier) for collaborative documentation generation.
+Both systems demonstrate significant improvements over baseline approaches but struggle with scaling to large repositories and maintaining cross-file consistency.
+Evaluation methodologies have evolved from traditional NLP metrics like BLEU [35] and ROUGE [23] to more sophisticated approaches.
+Roy et al.
+[41] and Mastropaolo et al.
+[29] questioned the effectiveness of automatic metrics, while recent work has adopted LLM-as-a-judge methodologies [56, 25].
+Multi-dimensional evaluation frameworks now assess completeness, helpfulness, and truthfulness, recognizing that documentation quality cannot be reduced to a single dimension.
+2.3 Multi-Agent Systems for Software Engineering
+Multi-agent architectures have gained prominence for tackling complex software engineering tasks requiring diverse expertise and collaborative problem-solving.
+AutoGen [51] introduced a framework enabling multi-agent communication with language models, allowing agents to assume different roles and collaborate on complex tasks.
+MetaGPT [18] demonstrated meta-programming for collaborative frameworks, where agents follow software development methodologies to produce comprehensive solutions.
+Multi-agent approaches include HyperAgent [36], which presents generalist software engineering agents capable of solving coding tasks at scale, and AgileCoder [33], which introduces dynamic collaborative agents for software development based on agile methodology.
+In the code generation domain, MapCoder [55] employed map-reduce style collaboration to break down complex programming problems, while ChatDev [38] showed that multi-agent systems can manage the entire software development lifecycle.
+The integration of multi-agent systems with documentation represents a recent innovation, with DocAgent demonstrating how role-based collaboration can improve documentation quality through specialized agents for analysis, information retrieval, drafting, and verification.
+However, existing approaches often lack systematic dependency management and struggle with maintaining consistency across large codebases.
 3 Our Approach - CodeWiki
@@
 Leaf nodes represent specific requirements assessed by Judge Agent, while parent scores are computed through weighted aggregation.
-The hierarchical organization ensures comprehensive coverage from high-level architectural components down to specific implementation details, with weights reflecting the relative importance of each component for understanding the repository
+The hierarchical organization ensures comprehensive coverage from high-level architectural components down to specific implementation details.
 Following [44], we synthesize evaluation rubrics tailored for each repository, illustrated in Figure 2.
@@
 This demonstrates CodeWiki’s ability to synthesize both textual explanations and visual artifacts that capture architectural patterns and module relationships, enabling developers to quickly understand the system’s structure and design decisions.
-Appendix B Background and Related Work
-B.1 Code Summarization and Documentation
-B.1.1 Existing Approaches
-Advances in natural language processing (NLP) have enabled significant progress in the task of code summarization.
-Existing methods [13, 21, 18, 16, 1, 4, 17, 29] primarily focus on function-level documentation, and can broadly be divided into two categories based on the type of context they exploit (see Table 2 for a comprehensive comparison):
-Local-Level Context
-Early neural approaches established the foundation for automated documentation.
-CodeBERT [9] demonstrated effective bimodal training on code and natural language, while more recent LLM applications have shown substantial improvements.
-For instance, [13] reported 20.6 BLEU scores across six languages using GPT-3, an 11.2% improvement over prior methods.
-DocuMint [21] further explored the use of small language models for Python docstring generation with promising results.
-However, these approaches primarily rely on local code context and do not leverage repository-level information to improve documentation quality.
-Codebase-Level Context
-Recognizing the limitations of function-only context, recent work has explored broader codebase information to enrich documentation.
-[18] demonstrated that incorporating class- and repository-level context can substantially improve summary quality through few-shot learning and retrieval-augmented generation.
-ProConSuL [16] incorporated project-level information via compiler analysis and call graphs, employing a two-phase training procedure to reduce hallucinations.
-RepoAgent [17] introduced comprehensive LLM-powered repository documentation, though it was limited in content diversity and architectural depth.
-DocAgent [29] advanced the field through multi-agent collaboration with topological processing, employing specialized agents that respect code dependencies to improve completeness and accuracy.
-$\hookrightarrow$ Limitations and Our Motivation: Despite this progress, most approaches still target isolated snippets (e.g., functions or classes) and thus fail to capture the underlying architecture of the repository, as highlighted in Table 2.
-For a developer onboarding a completely new codebase, the primary challenge is not understanding what each individual function or class does, but rather grasping the high-level overview of the system, its dataflow and how different components interact.
-This underscores the need for an automatic methodology capable of generating documentation that spans from low-level details to high-level repository-wide perspectives, covering critical features and providing visual representations that support developer comprehension.
-Our approach, CodeWiki, addresses these limitations by generating true repository-level documentation while leveraging full codebase context and supporting multiple programming languages.
-B.1.2 Existing Evaluation Methods
-Most existing works evaluate the quality of generated documentation using textual similarity metrics such as BLEU or ROUGE.
-These metrics measure the overlap between generated text and reference ground-truth documentation in terms of lexical or token-level similarity.
-$\hookrightarrow$ Limitations and Our Motivation: While these methods provide a lightweight quantitative measure, they often fail to capture deeper semantic aspects of the documentation, such as clarity, correctness, and completeness in explaining the underlying code behavior.
-Recent studies [23, 8] have shown that traditional metrics like BLEU and ROUGE do not consistently correlate with human judgment of documentation quality.
-This gap highlights the need for an assessment system that can fairly and accurately evaluate documentation, ensuring that the generated content truly supports developers in comprehending the system.
-Table 2: Comparison of different code summarization approaches.
-Repo-Level Doc.
-indicates whether the approach generates repository-level documentation output.
-Codebase Context specifies whether the codebase information is used as input to generate the documentation.
-Approach Repo-Level Doc. Codebase Context Multilingual DocuMint ASAP ProConSul DocAgent RepoAgent CodeWiki
-B.2 Rubric-based evaluation
-A repository can be described through multiple representations, each using different structures and presentation styles while sharing the same underlying architecture.
-This diversity makes it difficult to evaluate open-ended outputs automatically, as no single metric can reliably evaluate which version is better.
-Although human evaluation provides trustworthy results, it is resource-intensive and slow.
-To overcome these limitations and facilitate rapid assessment, many benchmarks [26, 14, 24, 2] often employ large language models (LLMs) as automated evaluators, using task-specific rubrics to judge the quality of generated outputs.
-In our CodeWikiBench benchmark, we similarly construct a set of hierarchical rubrics based on the official documentation, providing a structured and fine-grained framework to evaluate the quality and completeness of generated repository-level documentation.
+Appendix B Related Work
+B.1 Code Intelligence and LLMs
+The rapid advancement of Large Language Models (LLMs) has revolutionized code intelligence tasks.
+Early work by von der Mosel et al.
+[47] demonstrated the effectiveness of Transformer-based models for software engineering, leading to specialized code models like CodeBERT [12], GraphCodeBERT [13], CodeT5 [49], CodeT5+ [50], and CodeTF [3].
+These established early transformer-based approaches for code understanding and generation.
+Comprehensive surveys by Niu et al.
+[34] and Hou et al.
+[19] established the importance of domain-specific pre-training, showing consistent improvements over general-purpose language models across code search, summarization, and generation tasks.
+However, recent studies highlight important limitations: Li et al.
+[22] found that models often make confident decisions without sufficient reasoning, while Zheng et al.
+[57] noted that LLM performance remains unstable and requires careful prompt engineering [42].
+For code generation, HumanEval [6] and MBPP [2] established foundational benchmarks, later extended to multiple languages by MultiPL-E [4] and CodeScope [52].
+More recently, CodeMMLU [32] has emerged as a comprehensive multi-task benchmark for assessing code understanding and reasoning capabilities of CodeLLMs.
+B.2 Automated Documentation Generation
+Documentation generation has emerged as a critical LLM application in software engineering.
+Traditional approaches focused on function-level summarization using datasets like CodeSearchNet [15] and FunCom [21].
+While Sun et al.
+[45] found that ChatGPT can generate coherent comments, their quality often falls short of specialized models when evaluated using BLEU and ROUGE metrics.
+Recent research has shifted toward repository-level documentation generation, which requires understanding complex dependencies and architectural context.
+Luo et al.
+[27] introduced RepoAgent, an LLM-powered framework that leverages static analysis and incremental processing.
+DocAgent [53] proposed a multi-agent system with specialized roles (Reader, Searcher, Writer, Verifier) for collaborative documentation generation.
+Both systems demonstrate significant improvements over baseline approaches but struggle with scaling to large repositories and maintaining cross-file consistency.
+Evaluation methodologies have evolved from traditional NLP metrics like BLEU [35] and ROUGE [23] to more sophisticated approaches.
+Roy et al.
+[41] and Mastropaolo et al.
+[29] questioned the effectiveness of automatic metrics, while recent work has adopted LLM-as-a-judge methodologies [56, 25].
+Multi-dimensional evaluation frameworks now assess completeness, helpfulness, and truthfulness, recognizing that documentation quality cannot be reduced to a single dimension.
+B.3 Multi-Agent Systems for Software Engineering
+Multi-agent architectures have gained prominence for tackling complex software engineering tasks requiring diverse expertise and collaborative problem-solving.
+AutoGen [51] introduced a framework enabling multi-agent communication with language models, allowing agents to assume different roles and collaborate on complex tasks.
+MetaGPT [18] demonstrated meta-programming for collaborative frameworks, where agents follow software development methodologies to produce comprehensive solutions.
+Multi-agent approaches include HyperAgent [36], which presents generalist software engineering agents capable of solving coding tasks at scale, and AgileCoder [33], which introduces dynamic collaborative agents for software development based on agile methodology.
+In the code generation domain, MapCoder [55] employed map-reduce style collaboration to break down complex programming problems, while ChatDev [38] showed that multi-agent systems can manage the entire software development lifecycle.
+The integration of multi-agent systems with documentation represents a recent innovation, with DocAgent demonstrating how role-based collaboration can improve documentation quality through specialized agents for analysis, information retrieval, drafting, and verification.
+However, existing approaches often lack systematic dependency management and struggle with maintaining consistency across large codebases.
 Appendix C Rubric Reliability Assessment
@@
 Table 3 presents comprehensive results using CodeWiki-kimi-k2 (documentation generated using Kimi K2 Instruct) compared against the DeepWiki baseline.
-Repository Language LOC System Score (%) Coverage Improvement (%)
+Repository Lang. LOC System Score (%) Cov. Imp. (%)
 chartjs–Chart.js JavaScript 90,925 DeepWiki 88.71 ± 3.29 57/61

### v4

@@
 Furthermore, the hierarchical nature of software systems requires documentation that can serve multiple audiences, from high-level stakeholders seeking architectural overviews to developers requiring detailed implementation guidance [31].
-$\hookrightarrow$ To address these challenges, we present CodeWiki, the first open-source framework specifically designed to generate holistic, structured repository-level documentation that includes diverse content types such as system architecture diagrams, data flow visualizations, and sequence diagrams while scaling effectively for projects of arbitrary size.
+$\hookrightarrow$ To address these challenges, we present CodeWiki, an open-source framework specifically designed to generate holistic, structured repository-level documentation that includes diverse content types such as system architecture diagrams, data flow visualizations, and sequence diagrams while scaling effectively for projects of arbitrary size.
 CodeWiki introduces several key innovations: a hierarchical decomposition strategy inspired by dynamic programming principles that breaks complex repositories into manageable modules while preserving architectural coherence, a recursive agentic system with dynamic delegation capabilities that enables adaptive processing based on module complexity, and a comprehensive documentation synthesis approach that generates multiple forms of visual and textual documentation reflecting fine-grained codebase collaboration patterns.
@@
 Repository-level evaluation is particularly challenging as documentation admits multiple valid structures sharing underlying architectural content, highlighting the need for robust assessment beyond surface-level similarity.
-$\hookrightarrow$ To fill this gap, we introduce CodeWikiBench, the first benchmark specifically designed for repository-level documentation.
+$\hookrightarrow$ To fill this gap, we introduce CodeWikiBench, a comprehensive benchmark specifically designed for repository-level documentation.
 Instead of relying on surface-level similarity, we derive hierarchical rubrics from the official documentation maintained by project developers, enabling the evaluation of underlying repository properties.
@@
   1. 1.
-    CodeWiki, the first open-source semi-agentic framework for scalable repository-level documentation with proprietary and open-source LLM backends.
+    CodeWiki, an open-source semi-agentic framework for scalable repository-level documentation with proprietary and open-source LLM backends.
   2. 2.
-    CodeWikiBench, the first repository-level documentation benchmark with hierarchical rubrics and agentic assessment.
+    CodeWikiBench, a comprehensive repository-level documentation benchmark with hierarchical rubrics and agentic assessment.
   3. 3.
-    First multilingual repository documentation supporting 7 widely used languages.
+    Comprehensive multilingual repository documentation supporting 7 widely used languages.
   4. 4.
@@
 Synthesis involves multiple stages: analyzing child documentation for themes and patterns, generating architectural overviews explaining module collaboration, creating feature summaries distilling capabilities, developing usage guides for public interfaces, and producing architectural diagrams visualizing relationships and data flows.
-4 The First Benchmarking System - CodeWikiBench
+4 Benchmarking System - CodeWikiBench
 Repository-level documentation evaluation presents unique challenges due to complexity and hierarchical organization.
@@
 5.2 Baseline
-To the best of our knowledge, CodeWiki represents the first open-source framework specifically designed for comprehensive repository-level documentation generation.
-The absence of comparable open-source solutions necessitates comparison with closed-source alternatives.
+CodeWiki provides an open-source framework specifically designed for comprehensive repository-level documentation generation.
+Due to the limited availability of comparable open-source solutions, we compare with closed-source alternatives.
 We select DeepWiki [10] as our primary baseline, as it shares our goal of automated repository-level documentation generation and has demonstrated effectiveness in industrial applications.
@@
 Additional evaluation on 14 diverse repositories in Appendix D further validates these findings, with CodeWiki-kimi-k2 achieving an average improvement of 1.25% and substantial gains on large-scale projects including Trino (+29.53%) and Storybook (+11.42%).
-Answer to RQ1: CodeWiki demonstrates superior overall documentation quality and coverage compared to existing systems.
+Answer to RQ1: CodeWiki demonstrates strong overall documentation quality and coverage compared to existing systems.
 With proprietary models, it achieves 4.73% improvement; with open-source models, it matches baseline performance while offering transparency and accessibility benefits.
@@
 CodeWiki-sonnet-4 shows strong improvements in high-level scripting languages (+10.47%) and managed/enterprise languages (+4.04%), while systems programming languages present challenges for both systems.
-CodeWiki-sonnet-4 demonstrates exceptional performance on high-level scripting languages, achieving 79.14% average across Python, JavaScript, and TypeScript repositories—a 10.47 percentage point improvement over DeepWiki’s 68.67%.
+CodeWiki-sonnet-4 demonstrates strong performance on high-level scripting languages, achieving 79.14% average across Python, JavaScript, and TypeScript repositories—a 10.47 percentage point improvement over DeepWiki’s 68.67%.
 CodeWiki-kimi-k2 achieves 77.31% on these languages, showing strong performance with open-source models (+8.64% over DeepWiki).
@@
 This suggests that specialized parsing strategies may benefit all approaches for handling complex constructs like pointer operations and template metaprogramming.
-Answer to RQ2: CodeWiki demonstrates strong cross-language generalization with exceptional performance on high-level languages (79.14% with proprietary models, 77.31% with open-source models) and consistent improvements on managed languages, while systems programming languages present similar challenges across all frameworks and model types.
+Answer to RQ2: CodeWiki demonstrates strong cross-language generalization with high performance on high-level languages (79.14% with proprietary models, 77.31% with open-source models) and consistent improvements on managed languages, while systems programming languages present similar challenges across all frameworks and model types.
 5.4.3 RQ3: Scalability and Reliability
@@
 These complementary results reinforce our findings from the main experiments, showing an average improvement of 1.25% across 14 diverse repositories.
-Notably, CodeWiki achieves substantial gains in several large-scale projects, including a remarkable 29.53% improvement on the Trino database engine (1.5M LOC) and 11.42% on the Storybook UI development platform (365K LOC).
+Notably, CodeWiki achieves substantial gains in several large-scale projects, including a 29.53% improvement on the Trino database engine (1.5M LOC) and 11.42% on the Storybook UI development platform (365K LOC).
 The results further validate that CodeWiki’s hierarchical approach generalizes well across different repository characteristics and programming paradigms.

### v5

@@
 Abstract
-Given a large and evolving codebase, the ability to automatically generate a holistic, architecture-aware documentation that captures not only individual functions but also their cross-file, cross-module, and system-level interactions remains an open challenge.
-Comprehensive documentation is essential for long-term software maintenance and team collaboration, yet current automated approaches still fail to model the rich semantic dependencies and architectural structures that define real-world software systems.
-We present CodeWiki, a unified framework for automated repository-level documentation across seven programming languages.
-CodeWiki introduces three key innovations: (i) hierarchical decomposition that preserves architectural context across multiple levels of granularity, (ii) recursive multi-agent processing with dynamic task delegation for scalable generation, and (iii) multi-modal synthesis that integrates textual descriptions with visual artifacts such as architecture diagrams and data-flow representations.
-To enable rigorous evaluation, we introduce CodeWikiBench, a comprehensive benchmark featuring multi-dimensional rubrics and LLM-based assessment protocols.
-Experimental results demonstrate that CodeWiki achieves a 68.79% quality score with proprietary models, outperforming the closed-source DeepWiki baseline (64.06%) by 4.73%, with particularly strong improvements on high-level scripting languages (+10.47%).
-We open-source CodeWiki to foster future research and community adoption.
-[…]
-[…]
+Comprehensive software documentation is crucial yet costly to produce.
+Despite recent advances in large language models (LLMs), generating holistic, architecture-aware documentation at the repository level remains challenging due to complex and evolving codebases that exceed LLM context limits.
+Existing automated methods struggle to capture rich semantic dependencies and architectural structure.
+We present CodeWiki, a unified framework for automated repository-level documentation across seven mainstream programming languages.
+CodeWiki combines top-down hierarchical decomposition with a divide-and-conquer agent system to preserve architectural context and scale documentation generation, and a bottom-up synthesis that integrates textual descriptions with visual artifacts such as architecture and data-flow diagrams.
+We also introduce CodeWikiBench, a benchmark with hierarchical rubrics and LLM-based evaluation protocols.
+Experiments show that CodeWiki achieves a 68.79% quality score with proprietary models, outperforming the closed-source DeepWiki baseline by 4.73%, with especially strong gains on scripting languages.
+CodeWiki is released as open source to support future research.
 1 Introduction
@@
 In the rapidly evolving landscape of software development, maintaining comprehensive and up-to-date documentation has become increasingly challenging as codebases grow in size and complexity.
-Recent industry surveys reveal that documentation represents a critical bottleneck in modern development workflows: approximately 31% of developers currently rely heavily on AI for documenting code, while 25% use AI extensively for creating and maintaining documentation—making documentation tasks among the highest priorities for AI integration over the next 3-5 years [43].
-Notably, while developers express high satisfaction with AI tools for searching and learning, they indicate plans to substantially increase AI adoption for documentation and testing tasks, suggesting current tools remain insufficient for these critical needs.
-This widespread focus on AI-driven documentation reflects its dual burden: while essential for software maintenance and collaboration [9, 58, 5], manual creation and maintenance remain labor-intensive and often neglected [30].
-The emergence of Large Language Models (LLMs) [54, 17] has opened opportunities for automating documentation generation.
+Studies indicate that developers spend approximately 58% of their working time on program comprehension activities (Xia et al., 2018), highlighting the critical importance of accessible and accurate documentation.
+Despite documentation being recognized as essential for software maintenance and collaboration (de Souza et al., 2005; Zhi et al., 2015; Chen and Huang, 2009), the manual creation and maintenance of repository-level documentation remain a labor-intensive and often neglected aspect of software engineering practice (McBurney et al., 2018).
+The emergence of Large Language Models (LLMs)  (Vaswani et al., 2017; Zhang et al., 2023; He et al., 2025b) has opened opportunities for automating documentation generation tasks.
 However, current approaches still face two fundamental limitations.
-Limitations of Scalability at the Repository Level
-Recent advances in code understanding and natural language generation have demonstrated promising results in function-level and file-level documentation tasks [12, 20, 37, 28, 26, 1, 7, 27, 53] However, existing approaches face significant challenges when scaling to repository-level documentation.
-Unlike function-level documentation that focuses on individual components, repository-level documentation task is to capture architectural patterns, cross-module interactions, data flows, and system-wide design decisions [39, 46].
-Furthermore, the hierarchical nature of software systems requires documentation that can serve multiple audiences, from high-level stakeholders seeking architectural overviews to developers requiring detailed implementation guidance [31].
-$\hookrightarrow$ To address these challenges, we present CodeWiki, an open-source framework specifically designed to generate holistic, structured repository-level documentation that includes diverse content types such as system architecture diagrams, data flow visualizations, and sequence diagrams while scaling effectively for projects of arbitrary size.
-CodeWiki introduces several key innovations: a hierarchical decomposition strategy inspired by dynamic programming principles that breaks complex repositories into manageable modules while preserving architectural coherence, a recursive agentic system with dynamic delegation capabilities that enables adaptive processing based on module complexity, and a comprehensive documentation synthesis approach that generates multiple forms of visual and textual documentation reflecting fine-grained codebase collaboration patterns.
-Limitations of the Evaluation Landscape
-Despite progress, systematic benchmarks remain lacking.
-Function-level and file-level work relies on general-purpose metrics (BLEU, ROUGE) that fail to capture documentation quality nuances [40, 11].
-Repository-level evaluation is particularly challenging as documentation admits multiple valid structures sharing underlying architectural content, highlighting the need for robust assessment beyond surface-level similarity.
-$\hookrightarrow$ To fill this gap, we introduce CodeWikiBench, a comprehensive benchmark specifically designed for repository-level documentation.
-Instead of relying on surface-level similarity, we derive hierarchical rubrics from the official documentation maintained by project developers, enabling the evaluation of underlying repository properties.
-For robustness, CodeWikiBench further adopts an agentic assessment framework in which judge agents evaluate generated documentation against rubric requirements and aggregate results hierarchically, producing both quality scores and reliability measures.
-Limitations of Multilingual Support
+Limitations of Scalability at the Repository Level.
+Recent advances in code understanding and natural language generation have demonstrated promising results in function and file documentation tasks (Feng et al., 2020; Khan and Uddin, 2023; Poudel et al., 2024; Makharev and Ivanov, 2025; Lomshakov et al., 2024; Ahmed and Devanbu, 2023; Choi et al., 2023; Luo et al., 2024; Yang et al., 2025).
+However, existing approaches face significant challenges when scaling to repository documentation.
+Unlike function documentation that focuses on individual components, repository documentation must capture architectural patterns, interactions between modules, data flows, and design decisions that span the entire system (Rai et al., 2022; Treude et al., 2020).
+Furthermore, the complex hierarchical nature of software systems requires documentation that can serve multiple purposes, such as for stakeholders seeking architectural overviews or developers requiring detailed implementation guidance (Nassif and Robillard, 2025).
+$\hookrightarrow$ To address these challenges, we present CodeWiki, an open-source framework designed to generate holistic repository documentation that includes diverse content types such as system architecture diagrams, data flow visualizations, and sequence diagrams while scaling effectively for projects of arbitrary size.
+While prior work such as RepoAgent (Luo et al., 2024) aggregates function and class documentation into repository structures, and open-source alternatives like Open-DeepWiki apply LLMs to entire repositories, CodeWiki differs through its hierarchical synthesis approach where parent modules are synthesized from child documentation, combined with integrated visual artifacts capturing architectural patterns.
+CodeWiki introduces three key innovations: (1) a hierarchical decomposition strategy inspired by dynamic programming principles that breaks complex repositories into manageable modules while preserving architectural coherence; (2) a scalable divide-and-conquer agent system that dynamically determines the relevant context for each module, recursively partitions it into sub-modules handled by dedicated sub-agents; and (3) a bottom-up synthesis process that integrates textual descriptions with visual artifacts such as architecture diagrams and data-flow representations to generate high-level documentation.
+Limitations of the Evaluation Landscape.
+Despite progress in automated documentation, there is still a lack of systematic and standardized benchmarks.
+Existing work at the function-level and file-level has mostly been evaluated using general-purpose text generation metrics (e.g., BLEU, ROUGE), which fail to capture the nuanced quality of software documentation (Roy et al., 2021; Evtikhiev et al., 2023).
+For repository-level documentation, the large scale and complexity of software systems make evaluation particularly challenging, as the quality of documentation cannot be adequately judged by surface-level text-similarity metrics alone.
+A single repository can be described through documentation with different structures and presentations, yet all share the same underlying architectural content.
+This highlights the lack of a robust evaluation system that can fairly assess documentation quality regardless of how it is presented.
+$\hookrightarrow$ To fill this gap, we introduce CodeWikiBench, the first benchmark specifically designed for repository-level documentation.
+Instead of relying on surface-level similarity, we automatically derive hierarchical evaluation rubrics from the ground-truth official documentation maintained by project developers, enabling the automated evaluation of machine-generated documentation using the derived rubrics.
+This approach is inspired by recent advances in hierarchical rubric-based evaluation (Starace et al., 2025), which demonstrates the effectiveness of decomposing complex evaluation tasks into smaller, well-defined sub-tasks with clear grading criteria.
+For robustness, CodeWikiBench adopts an agentic assessment framework in which judge agents evaluate generated documentation against rubric requirements and aggregate results hierarchically, producing both quality scores and reliability measures.
+Limitations of Multilingual Support.
 Most studies on code documentation generation have primarily focused on Python, with limited consideration of other widely used programming languages such as Java, JavaScript, C, or C++.
@@
 In addition, CodeWikiBench provides a multilingual benchmark built from multiple repositories in the same set of languages, enabling comprehensive evaluation.
-Using CodeWikiBench, we demonstrate CodeWiki achieves 68.79% quality score with proprietary models, outperforming closed-source DeepWiki, while open-source alternatives achieve 64.80%, producing comprehensive documentation with architectural diagrams, usage patterns, and dependency visualizations (Appendix A).
-In summary, the main contributions of this work are:
+Using CodeWikiBench, we evaluate the effectiveness of CodeWiki against multiple repository-level documentation systems, including open-source baselines (OpenDeepWiki [^1] https://github.com/AIDotNet/OpenDeepWiki, deepwiki-open [^2] https://github.com/AsyncFuncAI/deepwiki-open) and the closed-source DeepWiki.
+The results show that CodeWiki achieves an overall quality score of 68.79%, outperforming all baselines with improvements of up to 18.54% over the strongest baseline, DeepWiki, on individual repositories.
+CodeWiki produces comprehensive documentation encompassing architectural diagrams, usage patterns, and cross-module dependency visualizations (see Appendix A for an example).
+In summary, our main contributions are:
   1. 1.
-    CodeWiki, an open-source semi-agentic framework for scalable repository-level documentation with proprietary and open-source LLM backends.
+    CodeWiki, an open-source framework for scalable repository documentation generation combining top-down and bottom-up analyses.
   2. 2.
-    CodeWikiBench, a comprehensive repository-level documentation benchmark with hierarchical rubrics and agentic assessment.
+    CodeWikiBench, a benchmark for repository documentation featuring hierarchical rubrics and an agentic assessment framework.
   3. 3.
-    Comprehensive multilingual repository documentation supporting 7 widely used languages.
+    Support for documentation generation across seven programming languages (Python, Java, JavaScript, TypeScript, C, C++, and C#).
   4. 4.
-    Competitive performance on real-world repositories, improving over DeepWiki with both proprietary (Claude Sonnet 4) and open-source (Kimi K2 Instruct) models.
-2 Related Work
-2.1 Code Intelligence and LLMs
-The rapid advancement of Large Language Models (LLMs) has revolutionized code intelligence tasks.
-Early work by von der Mosel et al.
-[47] demonstrated the effectiveness of Transformer-based models for software engineering, leading to specialized code models like CodeBERT [12], GraphCodeBERT [13], CodeT5 [49], CodeT5+ [50], and CodeTF [3].
-These established early transformer-based approaches for code understanding and generation.
-Comprehensive surveys by Niu et al.
-[34] and Hou et al.
-[19] established the importance of domain-specific pre-training, showing consistent improvements over general-purpose language models across code search, summarization, and generation tasks.
-However, recent studies highlight important limitations: Li et al.
-[22] found that models often make confident decisions without sufficient reasoning, while Zheng et al.
-[57] noted that LLM performance remains unstable and requires careful prompt engineering [42].
-For code generation, HumanEval [6] and MBPP [2] established foundational benchmarks, later extended to multiple languages by MultiPL-E [4] and CodeScope [52].
-More recently, CodeMMLU [32] has emerged as a comprehensive multi-task benchmark for assessing code understanding and reasoning capabilities of CodeLLMs.
-2.2 Automated Documentation Generation
-Documentation generation has emerged as a critical LLM application in software engineering.
-Traditional approaches focused on function-level summarization using datasets like CodeSearchNet [15] and FunCom [21].
-While Sun et al.
-[45] found that ChatGPT can generate coherent comments, their quality often falls short of specialized models when evaluated using BLEU and ROUGE metrics.
-Recent research has shifted toward repository-level documentation generation, which requires understanding complex dependencies and architectural context.
-Luo et al.
-[27] introduced RepoAgent, an LLM-powered framework that leverages static analysis and incremental processing.
-DocAgent [53] proposed a multi-agent system with specialized roles (Reader, Searcher, Writer, Verifier) for collaborative documentation generation.
-Both systems demonstrate significant improvements over baseline approaches but struggle with scaling to large repositories and maintaining cross-file consistency.
-Evaluation methodologies have evolved from traditional NLP metrics like BLEU [35] and ROUGE [23] to more sophisticated approaches.
-Roy et al.
-[41] and Mastropaolo et al.
-[29] questioned the effectiveness of automatic metrics, while recent work has adopted LLM-as-a-judge methodologies [56, 25].
-Multi-dimensional evaluation frameworks now assess completeness, helpfulness, and truthfulness, recognizing that documentation quality cannot be reduced to a single dimension.
-2.3 Multi-Agent Systems for Software Engineering
-Multi-agent architectures have gained prominence for tackling complex software engineering tasks requiring diverse expertise and collaborative problem-solving.
-AutoGen [51] introduced a framework enabling multi-agent communication with language models, allowing agents to assume different roles and collaborate on complex tasks.
-MetaGPT [18] demonstrated meta-programming for collaborative frameworks, where agents follow software development methodologies to produce comprehensive solutions.
-Multi-agent approaches include HyperAgent [36], which presents generalist software engineering agents capable of solving coding tasks at scale, and AgileCoder [33], which introduces dynamic collaborative agents for software development based on agile methodology.
-In the code generation domain, MapCoder [55] employed map-reduce style collaboration to break down complex programming problems, while ChatDev [38] showed that multi-agent systems can manage the entire software development lifecycle.
-The integration of multi-agent systems with documentation represents a recent innovation, with DocAgent demonstrating how role-based collaboration can improve documentation quality through specialized agents for analysis, information retrieval, drafting, and verification.
-However, existing approaches often lack systematic dependency management and struggle with maintaining consistency across large codebases.
+    Experimental validation demonstrating improvements over all baselines, with gains of up to 18.54% on individual repositories and 4.73% on average compared to the closed-source DeepWiki system.
+A detailed discussion of related work and positioning relative to prior approaches is provided in Appendix B.
 2 Our Approach - CodeWiki
@@
 Dependency Graph Construction
-Following DocAgent [53], we employ Tree-Sitter parsers to extract ASTs, systematically identifying functions, methods, classes, structs, modules, and interdependencies: function calls, class inheritance, attribute access, and module imports.
-We normalize these to a unified depends_on relation for cross-language generalization, creating directed graph $G=(V,E)$ where edge $(u,v)\in E$ indicates component $u$ depends on $v$ [24].
+Following DocAgent (Yang et al., 2025), we employ Tree-Sitter parsers to extract Abstract Syntax Trees (ASTs).
+We systematically identify functions, methods, classes, structs, modules, and interdependencies including: function calls (runtime dependencies), class inheritance (structural hierarchies), attribute access (data dependencies), and module imports (compilation dependencies).
+We normalize these to a unified depends_on relation for cross-language generalization.
+This creates a directed graph $G=(V,E)$ where vertices $V$ represent components and edge $(u,v)\in E$ indicates component $u$ depends on $v$.
+Understanding $v$ is prerequisite to comprehending $u$, fundamental to our hierarchical processing (Liu et al., 2023).
 Entry Point Identification and Decomposition
-Topological sorting identifies zero-in-degree components—independent entry points where users interact (main functions, API endpoints, CLI, public interfaces).
-High-level components are hierarchically decomposed into manageable modules through recursive partitioning considering component interdependencies and semantic coherence.
-For scalability, only component IDs serve as input, resulting in a feature-oriented module tree.
+Topological sorting identifies zero-in-degree components, independent entry points where users interact with the project (main functions, API endpoints, CLI, public interfaces).
+Components from these files represent high-level features and documentation starting points.
+High-level components are hierarchically decomposed into manageable modules through recursive partitioning.
+For scalability, only component IDs (relative paths) serve as input.
+Decomposition considers: component interdependencies (grouping related functionality) and semantic coherence (preserving logical boundaries).
+The result is a feature-oriented module tree where each node contains component lists and sub-modules.
 2.2 Recursive Documentation Generation
-CodeWiki’s core innovation is recursive agent-based processing, enabling arbitrary repository size handling while maintaining bounded complexity and architectural coherence.
+CodeWiki’s core innovation is recursive agent processing, enabling arbitrary repository size handling while maintaining bounded complexity and architectural coherence.
 Agent Architecture
-Each leaf module is assigned a specialized agent equipped with: (1) complete source code access, (2) full module tree for cross-module understanding, (3) documentation workspace tools (view, create, edit operations), and (4) dependency graph traversal for contextual exploration.
-The agent workflow: analyzes components to understand functionality and interfaces, explores context by traversing dependencies within and across modules, and generates comprehensive markdown documentation including descriptions, usage examples, API specifications, and architectural insights.
+Each leaf module is assigned a specialized agent equipped with: (1) complete source code access, (2) full module tree for cross-module understanding, (3) documentation workspace tools, and (4) dependency graph traversal for contextual exploration.
+The agent workflow: analyzes components to understand functionality and interfaces, explores context by traversing dependencies, and generates comprehensive markdown documentation including descriptions, usage examples, API specifications, and architectural insights.
 Dynamic Delegation
+Our key innovation is adaptive scalability through dynamic delegation.
+When module complexity exceeds single-pass capacity, agents delegate submodules to specialized subagents.
+Delegation criteria include code complexity metrics, semantic diversity, and context window utilization.
+The recursive process follows bottom-up processing (see Algorithm 1 in Appendix C).
+Upon delegation, agents provide submodule specifications, the module tree updates, and newly created leaves process recursively.
+This enables handling modules of any size while maintaining quality.
+Cross-Module Reference Management
+Maintaining coherence across boundaries requires sophisticated reference management.
+When agents encounter external components, our resolution system creates cross-references rather than duplicating content.
+A global registry tracks documented components, enabling quick identification and navigation between related components.
+2.3 Hierarchical Assembly and Documentation Synthesis
+After leaf documentation, hierarchical assembly synthesizes component-level details into architectural overviews through recursive parent module processing.
+Parent modules undergo LLM synthesis with carefully crafted prompting.
+LLMs receive: child module documentation, module tree structure, dependency information, and synthesis instructions for architectural patterns and feature interactions.
+Synthesis involves multiple stages: analyzing child documentation for themes and patterns, generating architectural overviews explaining module collaboration, creating feature summaries distilling capabilities, developing usage guides for public interfaces, and producing architectural diagrams visualizing relationships and data flows.
+3 The First Benchmarking System - CodeWikiBench
+Figure 2: Example hierarchical evaluation rubric for the RAGFlow repository.
+The rubric mirrors the project’s architectural structure with weighted requirements at multiple levels.
+Leaf nodes represent specific requirements assessed by Judge Agent, while parent scores are computed through weighted aggregation.
+The hierarchical organization ensures comprehensive coverage from high-level architectural components down to specific implementation details, with weights reflecting the relative importance of each component for understanding the repository
+Repository-level documentation evaluation presents unique challenges due to complexity and hierarchical organization.
+Unlike code generation assessable through automated metrics, documentation quality requires nuanced evaluation of completeness, accuracy, and coherence across abstraction levels.
+Recent research highlights traditional metrics’ limitations for code tasks (Evtikhiev et al., 2023; Haldar and Hockenmaier, 2024), while LLM-as-a-Judge paradigms offer new possibilities (Wang et al., 2025; Crupi et al., 2025; He et al., 2025a).
+We develop a comprehensive framework synthesizing repository-specific rubrics and employing agentic assessment.
+3.1 Hierarchical Rubric Generation
+Following (Starace et al., 2025), we synthesize evaluation rubrics tailored for each repository, illustrated in Figure 2.
+We collect official documentation from selected open-source projects and hierarchically parse into structured JSON format, leveraging directory structure and markdown syntax.
+A Rubric Generator Agent processes each structure with tool access to detailed contents, generating hierarchical rubrics mirroring repository functionality.
+To enhance reliability, we employ multiple independent generations using different model families.
+Final rubrics emerge from synthesizing these perspectives, reducing single-model biases.
+Analysis shows generated rubrics demonstrate high consistency: 73.65% semantic reliability and 70.84% structural reliability (Appendix D).
+3.2 Documentation Assessment
+Our evaluation employs specialized Judge Agents operating under carefully designed protocols.
+Judges receive generated documentation in structured JSON format with hidden detailed contents, plus repository-specific rubrics.
+The systematic methodology: analyzes documentation structure and content, searches comprehensively for requirement coverage, makes binary adequacy decisions, and provides concise reasoning.
+Critically, judges evaluate only leaf-level requirements, ensuring assessment grounds in specific concrete criteria rather than abstract concepts.
+For example, in Figure 2, agents assess "DeepDoc Visual Parser for complex layouts" rather than broader "Document Processing Engine."
+For enhanced reliability, multiple Judge Agents using different model families evaluate each requirement. Final leaf scores average across assessments, reducing individual model biases.
+3.3 Hierarchical Score Aggregation
+Final quality scores compute through bottom-up weighted aggregation respecting rubric hierarchy while tracking reliability via standard deviation propagation.
+Let $R$ be the rubric root, $C(n)$ denote node $n$’s children, and $w(n)$ denote assigned weight.
+Leaf Assessment and Reliability
+For leaf nodes $\ell$, multiple assessments $s_{1},s_{2},\ldots,s_{m}$ aggregate as:
+$S(\ell)=\bar{s}=\frac{1}{m}\sum_{i=1}^{m}s_{i}$
+Reliability quantified via standard deviation:
+$\sigma_{\ell}=\sqrt{\frac{1}{m-1}\sum_{i=1}^{m}(s_{i}-\bar{s})^{2}}$
+Lower deviations indicate higher judge consensus and assessment reliability.
+Hierarchical Propagation
+For internal nodes $n$ with children $C(n)=\{c_{1},\ldots,c_{k}\}$ having scores $S(c_{i})$, weights $w(c_{i})$, and deviations $\sigma_{c_{i}}$:
+$S(n)=\frac{\sum_{i=1}^{k}w(c_{i})\cdot S(c_{i})}{\sum_{i=1}^{k}w(c_{i})}$
+$\sigma_{n}=\frac{\sqrt{\sum_{i=1}^{k}w(c_{i})^{2}\sigma_{c_{i}}^{2}}}{\sum_{i=1}^{k}w(c_{i})}$
+This ensures uncertainty propagates appropriately, with weights determining score contribution and reliability influence.
+Final Assessment
+Final repository score $S(R)\in[0,1]$ with deviation $\sigma_{R}$ represents weighted aggregation propagating assessments and reliability upward.
+Lower $\sigma_{R}$ indicates higher methodology reliability.
+This dual-metric approach provides quality assessment with confidence bounds, enabling nuanced interpretation.
+Weights reflect component importance based on criticality and complexity while maintaining statistical rigor in uncertainty quantification.
+Repository Language LOC System Score (%) Coverage Improvement (%)
+All-Hands-AI–OpenHands Python 229,909 OpenDeepWiki 58.12 ± 3.21 42/67
+deepwiki-open 61.35 ± 2.98 45/67
+DeepWiki 73.04 ± 2.54 54/67
+CodeWiki 82.45 ± 2.65 59/67 +9.41
+sveltejs–svelte JavaScript 124,576 OpenDeepWiki 55.23 ± 3.85 61/96
+deepwiki-open 57.89 ± 3.62 64/96
+DeepWiki 68.51 ± 3.31 76/96
+CodeWiki 71.96 ± 3.73 80/96 +3.45
+puppeteer–puppeteer TypeScript 136,302 OpenDeepWiki 51.82 ± 4.15 48/82
+deepwiki-open 54.67 ± 3.94 51/82
+DeepWiki 64.46 ± 3.72 60/82
+CodeWiki 83.00 ± 3.37 74/82 +18.54
+Unity-Technologies–ml-agents C# 86,106 OpenDeepWiki 62.45 ± 4.28 32/46
+deepwiki-open 65.12 ± 4.05 34/46
+DeepWiki 74.80 ± 3.69 39/46
+CodeWiki 79.78 ± 5.02 42/46 +4.98
+elastic–logstash Java 117,485 OpenDeepWiki 41.25 ± 4.52 28/57
+deepwiki-open 44.18 ± 4.31 31/57
+DeepWiki 54.80 ± 4.10 38/57
+CodeWiki 57.90 ± 3.43 38/57 +3.10
+wazuh–wazuh C 1,446,730 OpenDeepWiki 32.56 ± 5.82 18/46
+deepwiki-open 35.89 ± 5.45 21/46
+DeepWiki 68.68 ± 4.74 39/46
+CodeWiki 64.17 ± 5.44 34/46 -4.51
+electron–electron C++ 184,234 OpenDeepWiki 28.45 ± 3.95 35/92
+deepwiki-open 31.22 ± 3.78 38/92
+DeepWiki 44.10 ± 3.12 54/92
+CodeWiki 42.30 ± 3.26 48/92 -1.80
+Average OpenDeepWiki 47.13 ± 4.25
+deepwiki-open 50.05 ± 4.02
+DeepWiki 64.06 ± 3.60
+CodeWiki 68.79 ± 3.84 +4.73
+Table 1: Documentation generation performance comparison across repositories.
+Scores represent weighted averages with standard deviations indicating assessment consensus.
+Coverage shows satisfied leaf criteria out of total requirements.
+OpenDeepWiki, deepwiki-open are open-source alternatives; DeepWiki is the closed-source baseline.
+4 Experiment
+We conduct comprehensive experiments to assess CodeWiki’s effectiveness in generating repository-level documentation across diverse programming languages and project types.
+Our evaluation is structured around 3 core research questions that address the fundamental challenges of automated repository documentation generation.
+4.1 Research Questions
+Our experimental evaluation is guided by the following research questions:
+RQ1: Documentation Quality and Coverage How does CodeWiki compare to existing state-of-the-art repository documentation systems in terms of overall documentation quality and requirement coverage?
+RQ2: Language Generalization Does CodeWiki demonstrate consistent performance across diverse programming languages, particularly comparing high-level scripting languages (Python, JavaScript, TypeScript) versus systems programming languages (C, C++, C#, Java)?
+RQ3: Scalability and Reliability How does CodeWiki’s hierarchical decomposition approach perform across repositories of varying sizes and complexity, and what is the reliability of our evaluation methodology?
+4.2 Baselines
+We compare CodeWiki against multiple repository-level documentation systems spanning both open-source and closed-source solutions:
+Open-Source Baselines
+OpenDeepWiki and deepwiki-open are open-source alternatives applying LLMs to entire repositories for documentation generation.
+These systems represent the current state of open-source repository-level documentation tools.
+Closed-Source Baseline
+DeepWiki (DeepWiki, 2025) is a commercial system has demonstrated effectiveness in industrial applications for automated repository-level documentation generation.
+Positioning Relative to Function-Level Approaches
+We note that function-level documentation systems such as RepoAgent (Luo et al., 2024) and DocAgent (Yang et al., 2025) generate N separate documents for N components, representing a fundamentally different documentation paradigm from CodeWiki’s hierarchically synthesized module-level outputs.
+Direct quantitative comparison would require forcing one approach into evaluation criteria it was not designed to satisfy, so we focus our comparison on systems targeting similar repository-level documentation goals.
+4.3 Experimental Setup
+Repository Selection
+We curate a benchmark of 7 open-source repositories spanning diverse programming languages, project scales, and application domains.
+Our selection prioritizes: (1) language diversity for cross-language generalization, (2) varying codebase sizes ranging from 86K to 1.4M LOC, (3) actively maintained projects with high-quality official documentation for rubric construction, and (4) coverage of multiple domains such as web frameworks, automation tools, and machine learning platforms.
+Table 1 presents the repository characteristics, including primary languages and codebase sizes.
+The benchmark covers JavaScript (svelte), TypeScript (puppeteer), C (wazuh), C++ (electron), C# (ml-agents), Java (logstash), and Python (OpenHands), ensuring comprehensive coverage across major programming paradigms.
+Temporal Separation for Data Leakage Prevention
+To ensure evaluation integrity, we use repository snapshots from August-September 2025, which postdate the knowledge cutoffs of all models used in our experiments (Claude Sonnet 4: March 2025).
+This temporal gap provides strong evidence against data leakage concerns.
+Detailed commit information is provided in Appendix E.
+Model Configuration
+Our experimental configuration employs state-of-the-art language models: Claude Sonnet 4 for documentation generation, multiple models (Claude Sonnet 4, Gemini 2.5 Pro, and Kimi K2 Instruct) for rubric generation to reduce single-model bias, and three distinct judge models (Gemini 2.5 Flash, GPT OSS 120B, and Kimi K2 Instruct) for independent assessments ensuring robust evaluation through multi-model consensus.
+Configuration details are in Appendix G.
+Evaluation Protocol
+For each repository, we execute the CodeWiki pipeline and generate comprehensive documentation including architectural overviews, component descriptions, and usage guides.
+The evaluation process follows our hierarchical rubric methodology: (1) generating repository-specific rubrics from official documentation, (2) performing agentic assessment of generated documentation, and (3) computing weighted scores with uncertainty quantification through standard deviation tracking.
+4.4 Results and Analysis
+4.4.1 RQ1: Documentation Quality and Coverage
+Table 1 presents comprehensive evaluation results addressing the fundamental question of documentation quality.
+CodeWiki achieves an average documentation quality score of 68.79%, representing a 4.73 percentage point improvement over the closed-source DeepWiki baseline (64.06%) and substantial improvements over open-source alternatives OpenDeepWiki (47.13%) and deepwiki-open (50.05%).
+This demonstrates the effectiveness of our hierarchical synthesis approach compared to both simpler open-source methods and commercial solutions.
+The coverage metrics reveal that CodeWiki consistently satisfies more leaf-level requirements than baselines in most cases.
+Notably, CodeWiki outperforms all baselines in 5 out of 7 repositories, with particularly strong improvements in TypeScript (83.00% vs.
+DeepWiki’s 64.46%), Python (82.45% vs.
+73.04%), and JavaScript (71.96% vs.
+68.51%).
+The open-source alternatives show notably degraded performance on more complex codebases, suggesting that simpler whole-repository prompting approaches do not scale effectively.
+Answer: CodeWiki demonstrates superior overall documentation quality and coverage compared to both open-source and closed-source systems, achieving improvements of 21.66% over OpenDeepWiki, 18.74% over deepwiki-open, and 4.73% over DeepWiki.
+4.4.2 RQ2: Language Generalization
+Our analysis reveals distinct performance patterns across programming language categories.
+CodeWiki demonstrates strong performance on scripting languages, achieving average scores of 79.14% (Python, JavaScript, TypeScript) repositories - a 10.47% improvement over DeepWiki’s 68.67% average for the same languages.
+For managed languages (C#, Java), CodeWiki achieves competitive performance with moderate but consistent improvements of 4.98% and 3.10% respectively, averaging 68.84% versus DeepWiki’s 64.80%.
+Systems programming languages (C, C++) present challenges where both frameworks exhibit similar patterns: DeepWiki achieves 56.39% while CodeWiki reaches 53.24%.
+This suggests that specialized parsing strategies may benefit both approaches for handling complex constructs like pointer and template metaprogramming.
+Answer: CodeWiki demonstrates strong generalization with notable performance on scripting languages (79.14% average) and consistent improvements on managed languages, while systems programming languages present similar challenges for both frameworks.
+4.4.3 RQ3: Scalability and Reliability
+Repository size analysis demonstrates CodeWiki’s scalability across diverse project scales.
+Our framework processes repositories ranging from 86K LOC to 1.4M LOC, with performance consistency observed within language categories regardless of repository size.
+For high-level languages, CodeWiki maintains strong performance across different scales: Unity ML-Agents (86K LOC) achieves 79.78%, while the larger OpenHands (230K LOC) reaches 82.45%.
+This consistency demonstrates that our hierarchical approach successfully addresses context limitation challenges that typically constrain documentation generation at scale.
+Critically, both CodeWiki and DeepWiki exhibit similar behavioral patterns across different repository sizes within the same programming language families.
+Systems programming languages present consistent challenges for both frameworks regardless of project scale, suggesting that the primary factor affecting documentation quality is language-specific parsing complexity rather than repository size.
+The dynamic delegation mechanism in CodeWiki automatically adapts to varying complexity levels, maintaining bounded processing requirements through hierarchical decomposition.
+Reliability analysis through standard deviation measurements validates our multi-model assessment methodology.
+The average standard deviations of 3.84% (CodeWiki) versus 3.60% (DeepWiki) indicate comparable assessment reliability while achieving superior overall quality scores.
+A pilot human study (Appendix F) provides additional validation, showing that human preferences align with automated evaluation results (CodeWiki preferred in 7/9 assessments across 3 participants and 3 repositories).
+Answer: CodeWiki’s hierarchical decomposition enables consistent scalability across repository sizes from 86K to 1.4M LOC, with performance variations primarily attributable to language-specific characteristics rather than scale limitations.
+Human evaluation validates alignment between our automated methodology and human judgment.
+5 Conclusion
+We present CodeWiki, an open-source framework for automated repository documentation using hierarchical decomposition and recursive processing.
+CodeWiki performs bottom-up hierarchical synthesis, generating parent modules from children while integrating visual artifacts.
+Across 7 multilanguage repositories, CodeWiki achieves 68.79% quality score, outperforming DeepWiki’s 64.06%.
+Performance variations correlate primarily with language characteristics rather than repository size, with scripting languages showing consistent improvements while systems languages face challenges for all frameworks.
+Future work includes improved parsing for systems languages, version-aware documentation aligned with code evolution, and using documentation as semantic bridges for tasks like code migration.
+We release CodeWiki as open source to advance research in hierarchical documentation generation.
+Limitations
+While CodeWiki demonstrates strong performance across multiple languages and repository scales, several limitations warrant discussion.
+Systems Programming Languages
+Our framework exhibits reduced effectiveness on systems programming languages (C, C++) compared to high-level languages, achieving 53.24% versus 79.14% average scores.
+This gap stems from the inherent complexity of low-level constructs such as pointer operations, manual memory management, and template metaprogramming.
+For systems languages these constructs are the architecture, making this a limitation of our current analysis phase rather than a secondary concern.
+Developing specialized parsing modules that leverage language-specific characteristics represents an important direction for future work.
+Rubric Generation
+The automatically generated rubrics, while showing 73.65% semantic and 70.84% structural reliability across model families (Appendix D), have not undergone comprehensive human verification.
+Some generated rubrics may be difficult to interpret, and future work should incorporate systematic human validation to ensure rubric quality and clarity.
+Evaluation Methodology
+Our evaluation methodology relies on LLM-based assessment with multi-model consensus, which may introduce model-specific biases despite our mitigation strategies.
+A pilot human study (Appendix F) with 3 participants across 3 repositories showed alignment between human preferences and automated scores (CodeWiki preferred in 7/9 assessments), providing preliminary validation.
+However, broader human evaluation with more participants and systematic rubric-based assessment would further strengthen the benchmark’s validity.
+The average standard deviation of 3.84% indicates reasonable consistency, yet edge cases may benefit from expert validation.
+Appendices
+Appendix A Example Generated Documentation
+Figure 3: Example repository-level documentation generated by CodeWiki for the All-Hands-AI–OpenHands repository.
+The documentation includes a comprehensive overview, architectural diagrams showing the modular event-driven architecture, and hierarchical navigation of core components including the agent system, implementations, and event processing modules.
+Figure 3 shows an example of repository-level documentation generated by CodeWiki for the All-Hands-AI–OpenHands repository.
+The documentation provides a comprehensive overview including the repository’s purpose, end-to-end architecture with visual diagrams, and hierarchical navigation of core components.
+This demonstrates CodeWiki’s ability to synthesize both textual explanations and visual artifacts that capture architectural patterns and module relationships, enabling developers to quickly understand the system’s structure and design decisions.
+Appendix B Background and Related Work
+B.1 Code Summarization and Documentation
+B.1.1 Existing Approaches
+Advances in natural language processing (NLP) have enabled significant progress in the task of code summarization.
+Existing methods  (Khan and Uddin, 2023; Poudel et al., 2024; Makharev and Ivanov, 2025; Lomshakov et al., 2024; Ahmed and Devanbu, 2023; Choi et al., 2023; Luo et al., 2024; Yang et al., 2025) primarily focus on function-level documentation, and can broadly be divided into two categories based on the type of context they exploit (see Table 2 for a comprehensive comparison):
+Local-Level Context
+Early neural approaches established the foundation for automated documentation.
+CodeBERT (Feng et al., 2020) demonstrated effective bimodal training on code and natural language, while more recent LLM applications have shown substantial improvements.
+For instance, (Khan and Uddin, 2023) reported 20.6 BLEU scores across six languages using GPT-3, an 11.2% improvement over prior methods.
+DocuMint (Poudel et al., 2024) further explored the use of small language models for Python docstring generation with promising results.
+However, these approaches primarily rely on local code context and do not leverage repository-level information to improve documentation quality.
+Codebase-Level Context
+Recognizing the limitations of function-only context, recent work has explored broader codebase information to enrich documentation.
+(Makharev and Ivanov, 2025) demonstrated that incorporating class- and repository-level context can substantially improve summary quality through few-shot learning and retrieval-augmented generation.
+ProConSuL (Lomshakov et al., 2024) incorporated project-level information via compiler analysis and call graphs, employing a two-phase training procedure to reduce hallucinations.
+RepoAgent (Luo et al., 2024) introduced comprehensive LLM-powered repository documentation, though it was limited in content diversity and architectural depth.
+DocAgent (Yang et al., 2025) advanced the field through multi-agent collaboration with topological processing, employing specialized agents that respect code dependencies to improve completeness and accuracy.
+$\hookrightarrow$ Limitations and Our Motivation: Despite this progress, most approaches still target isolated snippets (e.g., functions or classes) and thus fail to capture the underlying architecture of the repository, as highlighted in Table 2.
+For a developer onboarding a completely new codebase, the primary challenge is not understanding what each individual function or class does, but rather grasping the high-level overview of the system, its dataflow and how different components interact.
+This underscores the need for an automatic methodology capable of generating documentation that spans from low-level details to high-level repository-wide perspectives, covering critical features and providing visual representations that support developer comprehension.
+Our approach, CodeWiki, addresses these limitations by generating true repository-level documentation while leveraging full codebase context and supporting multiple programming languages.
+B.2 Repository-Level Documentation Systems
+Several systems have addressed repository-level documentation generation with varying approaches and scopes:
+Open-Source Alternatives
+OpenDeepWiki and deepwiki-open are community-developed tools that apply LLMs to entire repositories.
+However, as shown in Table 1, these approaches exhibit degraded performance on larger codebases (averaging 47.13% and 50.05% respectively), suggesting that simpler whole-repository prompting does not scale effectively to complex projects.
+RepoAgent (Luo et al., 2024) aggregates function and class-level documentation into repository structures using dependency analysis.
+While effective for component cataloging, it does not perform hierarchical synthesis where parent modules summarize and contextualize their children.
+codebase2tutorial [^3] https://code2tutorial.com/ generates tutorial-style documentation using whole-repository prompting, focusing on educational walkthroughs rather than comprehensive architectural documentation.
+LARCH (Koreeda et al., 2023) targets README generation specifically, addressing a narrower scope than full repository documentation.
+Agent-as-a-Judge (Zhuge et al., 2024) includes an OpenWiki feature that leverages existing READMEs and applies LLM assessment, representing a different methodology focused on evaluation rather than generation.
+CodeWiki’s Distinguishing Characteristics
+CodeWiki differs from these approaches through: (1) true hierarchical synthesis where parent modules are generated by summarizing and contextualizing child documentation, (2) integrated visual artifacts (architecture diagrams, data flow visualizations) generated alongside textual content, (3) dynamic delegation enabling scalable processing of arbitrarily large repositories, and (4) cross-module reference management maintaining documentation coherence across module boundaries.
+B.2.1 Existing Evaluation Methods
+Most existing works evaluate the quality of generated documentation using textual similarity metrics such as BLEU or ROUGE.
+These metrics measure the overlap between generated text and reference ground-truth documentation in terms of lexical or token-level similarity.
+$\hookrightarrow$ Limitations and Our Motivation: While these methods provide a lightweight quantitative measure, they often fail to capture deeper semantic aspects of the documentation, such as clarity, correctness, and completeness in explaining the underlying code behavior.
+Recent studies (Roy et al., 2021; Evtikhiev et al., 2023) have shown that traditional metrics like BLEU and ROUGE do not consistently correlate with human judgment of documentation quality.
+This gap highlights the need for an assessment system that can fairly and accurately evaluate documentation, ensuring that the generated content truly supports developers in comprehending the system.
+Table 2: Comparison of different code summarization approaches.
+Repo-Level Doc.
+indicates whether the approach generates repository-level documentation output.
+Codebase Context specifies whether the codebase information is used as input to generate the documentation.
+Approach Repo-Level Doc. Codebase Context Multilingual DocuMint ASAP ProConSul DocAgent RepoAgent CodeWiki
+B.3 Rubric-based evaluation
+A repository can be described through multiple representations, each using different structures and presentation styles while sharing the same underlying architecture.
+This diversity makes it difficult to evaluate open-ended outputs automatically, as no single metric can reliably evaluate which version is better.
+Although human evaluation provides trustworthy results, it is resource-intensive and slow.
+To overcome these limitations and facilitate rapid assessment, many benchmarks (Starace et al., 2025; Lin et al., 2025; Sirdeshmukh et al., 2025; Arora et al., 2025) often employ large language models (LLMs) as automated evaluators, using task-specific rubrics to judge the quality of generated outputs.
+In our CodeWikiBench benchmark, we similarly construct a set of hierarchical rubrics based on the official documentation, providing a structured and fine-grained framework to evaluate the quality and completeness of generated repository-level documentation.
+Appendix C Algorithm Details
 Input: Module tree $T$, dependency graph $G$
@@
 Algorithm 1 Recursive Module Documentation
-Our key innovation is adaptive scalability through dynamic delegation.
-When module complexity exceeds single-pass capacity, agents delegate sub-modules to specialized sub-agents.
-Delegation criteria: code complexity metrics (cyclomatic complexity, nesting depth), semantic diversity (functionally distinct subcomponents), and context window utilization.
-The recursive process follows bottom-up processing (Algorithm 1).
-Upon delegation, agents provide sub-module specifications, the module tree updates, and newly created leaves process recursively.
-This handles modules of any size while maintaining quality.
-Cross-Module Reference Management
-Maintaining coherence across boundaries requires sophisticated reference management.
-When agents encounter external components during traversal, our intelligent resolution system creates cross-references rather than duplicating content.
-A global registry tracks documented components and locations, enabling quick identification of documented components.
-This ensures conciseness while preserving completeness, creating interconnected documentation reflecting the actual codebase structure, enabling easy navigation between related components.
-3.3 Hierarchical Assembly and Documentation Synthesis
-After leaf documentation, hierarchical assembly synthesizes component-level details into architectural overviews through recursive parent module processing.
-Parent modules undergo LLM synthesis with carefully crafted prompting.
-LLMs receive: child module documentation, module tree structure, dependency information, and synthesis instructions for architectural patterns and feature interactions.
-Synthesis involves multiple stages: analyzing child documentation for themes and patterns, generating architectural overviews explaining module collaboration, creating feature summaries distilling capabilities, developing usage guides for public interfaces, and producing architectural diagrams visualizing relationships and data flows.
-4 Benchmarking System - CodeWikiBench
-Repository-level documentation evaluation presents unique challenges due to complexity and hierarchical organization.
-Unlike code generation assessable through automated metrics, documentation quality requires nuanced evaluation of completeness, accuracy, and coherence across abstraction levels.
-Recent research highlights traditional metrics’ limitations for code tasks [11, 14], while LLM-as-a-Judge paradigms offer new possibilities [48, 8, 16].
-We develop a comprehensive framework synthesizing repository-specific rubrics and employing agentic assessment.
-4.1 Hierarchical Rubric Generation
-Figure 2: Example hierarchical evaluation rubric for the RAGFlow repository.
-The rubric mirrors the project’s architectural structure with weighted requirements at multiple levels.
-Leaf nodes represent specific requirements assessed by Judge Agent, while parent scores are computed through weighted aggregation.
-The hierarchical organization ensures comprehensive coverage from high-level architectural components down to specific implementation details.
-Following [44], we synthesize evaluation rubrics tailored for each repository, illustrated in Figure 2.
-We collect official documentation from selected open-source projects and hierarchically parse into structured JSON format, leveraging directory structure and markdown syntax.
-This captures multi-level technical documentation from architectural overviews to implementation details.
-A Rubric Generator Agent processes each structure with tool access to detailed contents, generating hierarchical rubrics mirroring repository functionality and aligning with actual documentation standards.
-To enhance reliability, we employ multiple independent generations using different model families.
-Final rubrics emerge from synthesizing these perspectives, reducing single-model biases.
-Analysis shows generated rubrics demonstrate high consistency: 73.65% semantic reliability and 70.84% structural reliability (Appendix C).
-4.2 Documentation Assessment
-Our evaluation employs specialized Judge Agents operating under carefully designed protocols.
-Judges receive generated documentation in structured JSON format with hidden detailed contents, plus repository-specific rubrics.
-The systematic methodology: analyzes documentation structure and content, searches comprehensively for requirement coverage, makes binary adequacy decisions, and provides concise reasoning.
-Critically, judges evaluate only leaf-level requirements, ensuring assessment grounds in specific concrete criteria rather than abstract concepts.
-For example, in Figure 2, agents assess "DeepDoc Visual Parser for complex layouts" rather than broader "Document Processing Engine."
-For enhanced reliability, multiple Judge Agents using different model families evaluate each requirement. Final leaf scores average across assessments, reducing individual model biases.
-4.3 Hierarchical Score Aggregation
-Final quality scores compute through bottom-up weighted aggregation respecting rubric hierarchy while tracking reliability via standard deviation propagation.
-Let $R$ be the rubric root, $C(n)$ denote node $n$’s children, and $w(n)$ denote assigned weight.
-Leaf Assessment and Reliability
-For leaf nodes $\ell$, multiple assessments $s_{1},s_{2},\ldots,s_{m}$ aggregate as:
-$S(\ell)=\bar{s}=\frac{1}{m}\sum_{i=1}^{m}s_{i}$
-Reliability quantified via standard deviation:
-$\sigma_{\ell}=\sqrt{\frac{1}{m-1}\sum_{i=1}^{m}(s_{i}-\bar{s})^{2}}$
-Lower deviations indicate higher judge consensus and assessment reliability.
-Hierarchical Propagation
-For internal nodes $n$ with children $C(n)=\{c_{1},\ldots,c_{k}\}$ having scores $S(c_{i})$, weights $w(c_{i})$, and deviations $\sigma_{c_{i}}$:
-$S(n)=\frac{\sum_{i=1}^{k}w(c_{i})\cdot S(c_{i})}{\sum_{i=1}^{k}w(c_{i})},\quad\sigma_{n}=\frac{\sqrt{\sum_{i=1}^{k}w(c_{i})^{2}\sigma_{c_{i}}^{2}}}{\sum_{i=1}^{k}w(c_{i})}$
-This ensures uncertainty propagates appropriately, with weights determining score contribution and reliability influence.
-Final Assessment
-Final repository score $S(R)\in[0,1]$ with deviation $\sigma_{R}$ represents comprehensive weighted aggregation propagating assessments and reliability upward.
-Lower $\sigma_{R}$ indicates higher methodology reliability.
-This dual-metric approach provides quality assessment with confidence bounds, enabling nuanced interpretation.
-Weights reflect component importance based on criticality and complexity while maintaining statistical rigor in uncertainty quantification.
-Repository Language LOC System Score (%) Coverage Improvement (%)
-All-Hands-AI–OpenHands Python 229,909 DeepWiki 73.04 ± 2.54 54/67
-CodeWiki-kimi-k2 80.20 ± 3.26 60/67 +7.16
-CodeWiki-sonnet-4 82.45 ± 2.65 59/67 +9.41
-sveltejs–svelte JavaScript 124,576 DeepWiki 68.51 ± 3.31 76/96
-CodeWiki-kimi-k2 70.41 ± 3.39 80/96 +1.90
-CodeWiki-sonnet-4 71.96 ± 3.73 80/96 +3.45
-puppeteer–puppeteer TypeScript 136,302 DeepWiki 64.46 ± 3.72 60/82
-CodeWiki-kimi-k2 81.33 ± 3.33 67/82 +16.87
-CodeWiki-sonnet-4 83.00 ± 3.37 74/82 +18.54
-Unity-Technologies–ml-agents C# 86,106 DeepWiki 74.80 ± 3.69 39/46
-CodeWiki-kimi-k2 69.64 ± 5.40 40/46 -5.16
-CodeWiki-sonnet-4 79.78 ± 5.02 42/46 +4.98
-elastic–logstash Java 117,485 DeepWiki 54.80 ± 4.10 38/57
-CodeWiki-kimi-k2 56.89 ± 2.96 37/57 +2.09
-CodeWiki-sonnet-4 57.90 ± 3.43 38/57 +3.10
-wazuh–wazuh C 1,446,730 DeepWiki 68.68 ± 4.74 39/46
-CodeWiki-kimi-k2 58.15 ± 6.20 32/46 -10.53
-CodeWiki-sonnet-4 64.17 ± 5.44 34/46 -4.51
-electron–electron C++ 184,234 DeepWiki 44.10 ± 3.12 54/92
-CodeWiki-kimi-k2 36.99 ± 3.98 43/92 -7.11
-CodeWiki-sonnet-4 42.30 ± 3.26 48/92 -1.80
-Average DeepWiki 64.06 ± 3.60
-CodeWiki-kimi-k2 64.80 ± 4.07 +0.74
-CodeWiki-sonnet-4 68.79 ± 3.84 +4.73
-Table 1: Documentation generation performance comparison across diverse repositories.
-Scores represent weighted averages with standard deviations indicating assessment consensus.
-Coverage shows satisfied leaf criteria out of total requirements.
-Blue-highlighted rows show CodeWiki-sonnet-4 results, while yellow-highlighted rows show CodeWiki-kimi-k2 results.
-Improvement column shows relative gain over DeepWiki baseline.
-5 Experiment
-We conduct comprehensive experiments to evaluate CodeWiki’s effectiveness in generating repository-level documentation across diverse programming languages and project types.
-Our evaluation is structured around three core research questions that address the fundamental challenges of automated repository documentation generation.
-5.1 Research Questions
-Our experimental evaluation is guided by the following research questions:
-RQ1: Documentation Quality and Coverage How does CodeWiki compare to existing state-of-the-art repository documentation systems in terms of overall documentation quality and requirement coverage?
-RQ2: Cross-Language Generalization Does CodeWiki demonstrate consistent performance across diverse programming languages, particularly comparing high-level scripting languages (Python, JavaScript, TypeScript) versus systems programming languages (C, C++, C#, Java)?
-RQ3: Scalability and Reliability How does CodeWiki’s hierarchical decomposition approach perform across repositories of varying sizes and complexity, and what is the reliability of our evaluation methodology?
-5.2 Baseline
-CodeWiki provides an open-source framework specifically designed for comprehensive repository-level documentation generation.
-Due to the limited availability of comparable open-source solutions, we compare with closed-source alternatives.
-We select DeepWiki [10] as our primary baseline, as it shares our goal of automated repository-level documentation generation and has demonstrated effectiveness in industrial applications.
-5.3 Experimental Setup
-5.3.1 Repository Selection
-We carefully curate a comprehensive benchmark consisting of 21 open-source repositories that represent diverse programming languages, project scales, and application domains.
-The selection criteria prioritize:
-  1. 1.
-    Language diversity to test cross-language generalization
-  2. 2.
-    Varying codebase sizes from 16K to 1.7M lines of code
-  3. 3.
-    Active maintenance with comprehensive official documentation for rubric generation
-  4. 4.
-    Diverse application domains including web frameworks, automation tools, machine learning platforms, and system utilities
-Table 1 presents the main evaluation set of 7 repositories with detailed analysis, while Table 3 in Appendix D provides results for 14 additional repositories that further validate our approach across diverse codebases.
-5.3.2 Model Configuration
-We evaluate CodeWiki with two different LLM backends:
-  • CodeWiki-sonnet-4: Uses Claude Sonnet 4 (proprietary model)
-  • CodeWiki-kimi-k2: Uses Kimi K2 Instruct (open-source model)
-For rubric generation, we employ multiple models (Claude Sonnet 4, Gemini 2.5 Pro, Kimi K2 Instruct) to reduce single-model bias.
-The evaluation process uses three independent judge models (Gemini 2.5 Flash, GPT OSS 120B, Kimi K2 Instruct) for robust assessment.
-Additional implementation details are provided in Appendix E.
-5.3.3 Evaluation Protocol
-For each repository, we execute the complete CodeWiki pipeline and generate comprehensive documentation including architectural overviews, component descriptions, and usage guides.
-The evaluation process follows our hierarchical rubric methodology:
-  1. 1.
-    Rubric Generation: Repository-specific rubrics derived from official documentation
-  2. 2.
-    Agentic Assessment: Multiple judge agents evaluate generated documentation against rubric criteria
-  3. 3.
-    Score Aggregation: Weighted scores computed with uncertainty quantification through standard deviation tracking
-5.4 Results and Analysis
-5.4.1 RQ1: Documentation Quality and Coverage
-Table 1 presents comprehensive evaluation results addressing documentation quality.
-CodeWiki-sonnet-4 achieves an average quality score of 68.79%, representing a 4.73 percentage point improvement over DeepWiki’s 64.06%.
-Notably, CodeWiki-kimi-k2 demonstrates competitive performance at 64.80%, closely matching DeepWiki despite using an open-source model.
-This validates our framework’s effectiveness across different LLM backends.
-The coverage metrics reveal that CodeWiki consistently satisfies more leaf-level requirements than the baseline.
-CodeWiki-sonnet-4 outperforms DeepWiki in 5 out of 7 repositories, with particularly strong improvements in TypeScript (83.00% vs.
-64.46%, +18.54%), Python (82.45% vs.
-73.04%, +9.41%), and JavaScript (71.96% vs.
-68.51%, +3.45%).
-CodeWiki-kimi-k2 shows notable gains in TypeScript (+16.87%) and Python (+7.16%), demonstrating that our hierarchical approach enables strong performance even with open-source models.
-Additional evaluation on 14 diverse repositories in Appendix D further validates these findings, with CodeWiki-kimi-k2 achieving an average improvement of 1.25% and substantial gains on large-scale projects including Trino (+29.53%) and Storybook (+11.42%).
-Answer to RQ1: CodeWiki demonstrates strong overall documentation quality and coverage compared to existing systems.
-With proprietary models, it achieves 4.73% improvement; with open-source models, it matches baseline performance while offering transparency and accessibility benefits.
-5.4.2 RQ2: Cross-Language Generalization
-Our analysis reveals distinct performance patterns across programming language categories, as visualized in Figure 3.
-Figure 3: Performance comparison by programming language categories.
-CodeWiki-sonnet-4 shows strong improvements in high-level scripting languages (+10.47%) and managed/enterprise languages (+4.04%), while systems programming languages present challenges for both systems.
-CodeWiki-sonnet-4 demonstrates strong performance on high-level scripting languages, achieving 79.14% average across Python, JavaScript, and TypeScript repositories—a 10.47 percentage point improvement over DeepWiki’s 68.67%.
-CodeWiki-kimi-k2 achieves 77.31% on these languages, showing strong performance with open-source models (+8.64% over DeepWiki).
-For managed/enterprise languages (C#, Java), CodeWiki-sonnet-4 achieves competitive performance with an average improvement of 4.04% (68.84% vs.
-64.80%), while CodeWiki-kimi-k2 shows 63.27% (-1.53%).
-Systems programming languages present unique challenges where both frameworks exhibit similar patterns: CodeWiki-sonnet-4 achieves 53.24% compared to DeepWiki’s 56.39%, while CodeWiki-kimi-k2 obtains 47.58%.
-This suggests that specialized parsing strategies may benefit all approaches for handling complex constructs like pointer operations and template metaprogramming.
-Answer to RQ2: CodeWiki demonstrates strong cross-language generalization with high performance on high-level languages (79.14% with proprietary models, 77.31% with open-source models) and consistent improvements on managed languages, while systems programming languages present similar challenges across all frameworks and model types.
-5.4.3 RQ3: Scalability and Reliability
-Repository size analysis demonstrates CodeWiki’s scalability across 86K-1.4M LOC.
-For high-level languages, CodeWiki maintains strong performance across scales: Unity ML-Agents (86K LOC) achieves 79.78%, while OpenHands (230K LOC) reaches 82.45%, demonstrating that hierarchical decomposition successfully addresses context limitations.
-Both CodeWiki and DeepWiki exhibit similar behavioral patterns across repository sizes within language families, with performance variations primarily attributable to language-specific parsing complexity rather than scale.
-Dynamic delegation automatically adapts to complexity levels, maintaining bounded processing requirements.
-Reliability analysis through standard deviations (3.84% CodeWiki vs.
-3.60% DeepWiki) validates our multi-model assessment methodology.
-Lower deviations for high-level languages (Python: 2.65%, JavaScript: 3.73%) versus systems languages (C: 5.44%, C++: 3.26%) confirm evaluation consistency correlates with language characteristics rather than repository scale.
-Answer to RQ3: CodeWiki’s hierarchical decomposition enables consistent scalability across repository sizes from 86K to 1.4M LOC, with performance variations primarily attributable to language-specific characteristics rather than scale limitations.
-Both frameworks exhibit similar patterns across different repository sizes within language families.
-6 Conclusion
-We presented CodeWiki, an open-source semi-agentic framework addressing automated repository-level documentation through hierarchical decomposition and recursive processing.
-CodeWiki achieves 68.79% quality with proprietary models (Claude Sonnet 4) and 64.80% with open-source alternatives (Kimi K2 Instruct), both outperforming or matching DeepWiki’s 64.06%, demonstrating strong performance regardless of underlying model choice.
-Results show performance variations correlate primarily with language characteristics rather than repository size or model type.
-High-level languages show consistent improvements across scales and models, while systems languages face challenges regardless of configuration.
-Future directions include: developing specialized parsing modules for improved systems language support, extending to multi-version documentation tracking codebase evolution, and leveraging comprehensive documentation for downstream tasks like code migration.
-By releasing CodeWiki as open source with support for both proprietary and open-source LLMs, we accelerate research and enable the community to build upon our hierarchical processing approach with accessible technology choices.
-Limitations
-While CodeWiki demonstrates strong performance across multiple languages and repository scales, two main limitations warrant discussion.
-First, our framework exhibits reduced effectiveness on systems programming languages (C, C++) compared to high-level languages, achieving 53.24% versus 79.14% average scores.
-This gap stems from the inherent complexity of low-level constructs such as pointer operations, manual memory management, and template metaprogramming, which challenge both AST parsing and semantic understanding.
-Second, our evaluation methodology, while more comprehensive than traditional metrics, relies on LLM-based assessment which may introduce model-specific biases despite our multi-model consensus approach.
-The average standard deviation of 3.84% indicates reasonable consistency, yet edge cases may benefit from human expert validation.
-Appendices
-Appendix A Example Generated Documentation
-Figure 4: Example repository-level documentation generated by CodeWiki for the All-Hands-AI–OpenHands repository.
-The documentation includes a comprehensive overview, architectural diagrams showing the modular event-driven architecture, and hierarchical navigation of core components including the agent system, implementations, and event processing modules.
-Figure 4 shows an example of repository-level documentation generated by CodeWiki for the All-Hands-AI–OpenHands repository.
-The documentation provides a comprehensive overview including the repository’s purpose, end-to-end architecture with visual diagrams, and hierarchical navigation of core components.
-This demonstrates CodeWiki’s ability to synthesize both textual explanations and visual artifacts that capture architectural patterns and module relationships, enabling developers to quickly understand the system’s structure and design decisions.
-Appendix B Related Work
-B.1 Code Intelligence and LLMs
-The rapid advancement of Large Language Models (LLMs) has revolutionized code intelligence tasks.
-Early work by von der Mosel et al.
-[47] demonstrated the effectiveness of Transformer-based models for software engineering, leading to specialized code models like CodeBERT [12], GraphCodeBERT [13], CodeT5 [49], CodeT5+ [50], and CodeTF [3].
-These established early transformer-based approaches for code understanding and generation.
-Comprehensive surveys by Niu et al.
-[34] and Hou et al.
-[19] established the importance of domain-specific pre-training, showing consistent improvements over general-purpose language models across code search, summarization, and generation tasks.
-However, recent studies highlight important limitations: Li et al.
-[22] found that models often make confident decisions without sufficient reasoning, while Zheng et al.
-[57] noted that LLM performance remains unstable and requires careful prompt engineering [42].
-For code generation, HumanEval [6] and MBPP [2] established foundational benchmarks, later extended to multiple languages by MultiPL-E [4] and CodeScope [52].
-More recently, CodeMMLU [32] has emerged as a comprehensive multi-task benchmark for assessing code understanding and reasoning capabilities of CodeLLMs.
-B.2 Automated Documentation Generation
-Documentation generation has emerged as a critical LLM application in software engineering.
-Traditional approaches focused on function-level summarization using datasets like CodeSearchNet [15] and FunCom [21].
-While Sun et al.
-[45] found that ChatGPT can generate coherent comments, their quality often falls short of specialized models when evaluated using BLEU and ROUGE metrics.
-Recent research has shifted toward repository-level documentation generation, which requires understanding complex dependencies and architectural context.
-Luo et al.
-[27] introduced RepoAgent, an LLM-powered framework that leverages static analysis and incremental processing.
-DocAgent [53] proposed a multi-agent system with specialized roles (Reader, Searcher, Writer, Verifier) for collaborative documentation generation.
-Both systems demonstrate significant improvements over baseline approaches but struggle with scaling to large repositories and maintaining cross-file consistency.
-Evaluation methodologies have evolved from traditional NLP metrics like BLEU [35] and ROUGE [23] to more sophisticated approaches.
-Roy et al.
-[41] and Mastropaolo et al.
-[29] questioned the effectiveness of automatic metrics, while recent work has adopted LLM-as-a-judge methodologies [56, 25].
-Multi-dimensional evaluation frameworks now assess completeness, helpfulness, and truthfulness, recognizing that documentation quality cannot be reduced to a single dimension.
-B.3 Multi-Agent Systems for Software Engineering
-Multi-agent architectures have gained prominence for tackling complex software engineering tasks requiring diverse expertise and collaborative problem-solving.
-AutoGen [51] introduced a framework enabling multi-agent communication with language models, allowing agents to assume different roles and collaborate on complex tasks.
-MetaGPT [18] demonstrated meta-programming for collaborative frameworks, where agents follow software development methodologies to produce comprehensive solutions.
-Multi-agent approaches include HyperAgent [36], which presents generalist software engineering agents capable of solving coding tasks at scale, and AgileCoder [33], which introduces dynamic collaborative agents for software development based on agile methodology.
-In the code generation domain, MapCoder [55] employed map-reduce style collaboration to break down complex programming problems, while ChatDev [38] showed that multi-agent systems can manage the entire software development lifecycle.
-The integration of multi-agent systems with documentation represents a recent innovation, with DocAgent demonstrating how role-based collaboration can improve documentation quality through specialized agents for analysis, information retrieval, drafting, and verification.
-However, existing approaches often lack systematic dependency management and struggle with maintaining consistency across large codebases.
-Appendix C Rubric Reliability Assessment
+Appendix D Rubric Reliability Assessment
 To ensure the quality and trustworthiness of automatically generated evaluation rubrics, we developed a comprehensive reliability assessment framework that quantifies the consistency and structural integrity of rubrics produced by different language models.
@@
 Given a set of $n$ rubrics $\{R_{1},R_{2},\ldots,R_{n}\}$ produced by different models for the same documentation corpus, we compute pairwise consistency scores across both semantic and structural dimensions.
-C.1 Semantic Consistency Analysis
+D.1 Semantic Consistency Analysis
 We evaluate semantic consistency by extracting all requirement texts from each rubric and compute semantic similarity using distributed representations.
@@
 This formulation ensures that both the coverage of requirements from $R_{i}$ in $R_{j}$ and vice versa are considered, providing a balanced measure of semantic overlap.
-C.2 Structural Consistency Analysis
+D.2 Structural Consistency Analysis
 Structural consistency evaluates the architectural similarity between rubrics, focusing on hierarchical organization rather than content semantics.
@@
 $\quad+\text{Sim}_{\text{weights}}(R_{i},R_{j})\Big)$ (7)
-C.3 Overall Reliability Score
+D.3 Overall Reliability Score
 The overall reliability score aggregates consistency measures across all model pairs. For $n$ models, we compute $\binom{n}{2}$ pairwise comparisons and derive summary statistics:
@@
 Table 3: Rubric Reliability Assessment: Semantic and Structural Consistency Scores
-Repository Semantic (%) Structural (%)
-All-Hands-AI–OpenHands 73.64 68.82
-sveltejs–svelte 75.56 68.88
-puppeteer–puppeteer 72.97 76.36
-Unity-Technologies–ml-agents 75.95 71.22
-elastic–logstash 73.13 68.22
-wazuh–wazuh 72.20 67.62
-electron–electron 72.12 74.79
-Average 73.65 70.84
-Appendix D Complementary Experimental Results
-To further validate CodeWiki’s effectiveness across diverse programming languages and repository types, we conducted additional experiments on 14 repositories spanning 7 programming languages.
-Table 3 presents comprehensive results using CodeWiki-kimi-k2 (documentation generated using Kimi K2 Instruct) compared against the DeepWiki baseline.
-Repository Lang. LOC System Score (%) Cov. Imp. (%)
-chartjs–Chart.js JavaScript 90,925 DeepWiki 88.71 ± 3.29 57/61
-CodeWiki-kimi-k2 84.01 ± 3.55 54/61 -4.70
-marktext–marktext JavaScript 61,574 DeepWiki 81.64 ± 4.67 50/57
-CodeWiki-kimi-k2 82.49 ± 2.51 51/57 +0.85
-storybookjs–storybook TypeScript 364,606 DeepWiki 72.07 ± 3.59 64/78
-CodeWiki-kimi-k2 83.49 ± 2.31 69/78 +11.42
-mermaid-js–mermaid TypeScript 157,160 DeepWiki 73.81 ± 4.05 59/67
-CodeWiki-kimi-k2 71.52 ± 5.36 57/67 -2.29
-qmk–qmk_firmware C 1,099,824 DeepWiki 81.23 ± 2.62 94/110
-CodeWiki-kimi-k2 72.01 ± 2.63 88/110 -9.22
-sumatrapdfreader–sumatrapdf C 1,699,044 DeepWiki 55.81 ± 6.29 31/40
-CodeWiki-kimi-k2 58.92 ± 5.18 31/40 +3.11
-x64dbg–x64dbg C++ 207,671 DeepWiki 64.42 ± 2.07 62/107
-CodeWiki-kimi-k2 58.01 ± 3.58 56/107 -6.41
-nlohmann–json C++ 119,542 DeepWiki 66.06 ± 3.08 33/57
-CodeWiki-kimi-k2 61.28 ± 2.35 30/57 -4.78
-FluentValidation–FluentValidation C# 16,452 DeepWiki 76.48 ± 2.33 41/48
-CodeWiki-kimi-k2 70.03 ± 2.90 38/48 -6.45
-git-ecosystem–git-credential-manager C# 40,812 DeepWiki 87.03 ± 2.58 66/70
-CodeWiki-kimi-k2 82.78 ± 0.90 61/70 -4.25
-material-components–material-components-android Java 244,092 DeepWiki 70.99 ± 2.92 79/96
-CodeWiki-kimi-k2 79.16 ± 4.32 89/96 +8.17
-trinodb–trino Java 1,535,202 DeepWiki 55.27 ± 3.76 48/68
-CodeWiki-kimi-k2 84.80 ± 2.84 62/68 +29.53
-RasaHQ–rasa Python 147,711 DeepWiki 80.94 ± 3.15 64/70
-CodeWiki-kimi-k2 81.10 ± 2.69 64/70 +0.16
-microsoft–graphrag Python 30,390 DeepWiki 76.34 ± 3.28 48/56
-CodeWiki-kimi-k2 79.29 ± 2.96 50/56 +2.95
-Average DeepWiki 73.62 ± 3.41
-CodeWiki-kimi-k2 74.87 ± 3.15 +1.25
-Table 3: Complementary documentation generation performance evaluation across 14 additional repositories.
-Results demonstrate CodeWiki’s consistent effectiveness across diverse codebases, with notable improvements in Java repositories (trinodb–trino: +29.53%, material-components-android: +8.17%) and TypeScript (storybookjs–storybook: +11.42%).
-These complementary results reinforce our findings from the main experiments, showing an average improvement of 1.25% across 14 diverse repositories.
-Notably, CodeWiki achieves substantial gains in several large-scale projects, including a 29.53% improvement on the Trino database engine (1.5M LOC) and 11.42% on the Storybook UI development platform (365K LOC).
-The results further validate that CodeWiki’s hierarchical approach generalizes well across different repository characteristics and programming paradigms.
-Appendix E Implementation Details
-E.1 Model Configurations
+Repository Semantic (%) Structural (%) All-Hands-AI–OpenHands 73.64 68.82 sveltejs–svelte 75.56 68.88 puppeteer–puppeteer 72.97 76.36 Unity-Technologies–ml-agents 75.95 71.22 elastic–logstash 73.13 68.22 wazuh–wazuh 72.20 67.62 electron–electron 72.12 74.79 Average 73.65 70.84
+Appendix E Temporal Separation Details
+Table 4 presents the specific commit information for each benchmark repository, demonstrating temporal separation from model training data.
+Table 4: Repository commit information demonstrating temporal separation from model training data.
+Repository Language Commit ID Commit Date OpenHands Python 30604c4 Sep 13, 2025 svelte JavaScript be645b4 Aug 31, 2025 puppeteer TypeScript c1105f1 Aug 29, 2025 ml-agents C# 4cf2f49 Aug 15, 2025 logstash Java 895cfa5 Aug 29, 2025 wazuh C 44b7cd3 Aug 30, 2025 electron C++ 828fd59 Aug 30, 2025
+Appendix F Human Evaluation Pilot Study
+To provide additional validation of our automated evaluation methodology beyond multi-model consensus, we conducted a pilot human study comparing CodeWiki and DeepWiki documentation outputs.
+F.1 Study Design
+Participants
+We recruited three participants with diverse professional backgrounds relevant to documentation consumption:
+  • Business Analyst: Represents stakeholders seeking high-level architectural understanding
+  • AI Researcher: Represents technical users evaluating documentation for research purposes
+  • Software Engineer: Represents developers using documentation for implementation guidance
+None of the participants were prior contributors to the evaluated repositories, simulating the common scenario where developers approach unfamiliar codebases.
+Evaluation Protocol
+For each of three repositories (OpenHands, svelte, puppeteer), participants:
+  1. 1.
+    Received access to the official codebase and its official documentation
+  2. 2.
+    Reviewed anonymized documentation outputs from both CodeWiki and DeepWiki (presentation order randomized)
+  3. 3.
+    Assessed which documentation better reflected: (a) correctness relative to official documentation, and (b) coverage of actual repository architecture
+  4. 4.
+    Provided a binary preference judgment
+The evaluation was conducted blind, with system identities hidden from participants.
+F.2 Results
+Table 5: Pilot human evaluation results showing participant preferences between CodeWiki (CW) and DeepWiki (DW) across three repositories.
+Repository BA AI SE CW Pref.
+OpenHands CW CW CW 3/3
+svelte CW CW DW 2/3
+puppeteer CW DW CW 2/3
+Total 3/3 2/3 2/3 7/9 (77.8%)
+Table 5 presents the evaluation results. CodeWiki was preferred in 7 out of 9 assessments (77.8%). Key observations:
+Alignment with Automated Scores
+Human preferences correlate with the magnitude of automated evaluation improvements:
+  • OpenHands (+9.41% automated improvement): Unanimous CodeWiki preference (3/3)
+  • puppeteer (+18.54% automated improvement): Strong CodeWiki preference (2/3)
+  • svelte (+3.45% automated improvement): Closer preferences (2/3 for CodeWiki)
+This alignment between human judgment and automated evaluation supports the validity of our LLM-based assessment methodology.
+F.3 Limitations
+We acknowledge several limitations of this pilot study:
+  • Scale: Three participants and three repositories provide preliminary validation but not statistical significance
+  • Simplified Protocol: Due to budget constraints, we used preference judgment rather than full rubric-based evaluation
+  • Repository Selection: The evaluated repositories may not represent the full diversity of our benchmark
+Comprehensive human evaluation with larger participant pools, more repositories, and detailed rubric-based assessment represents important future work to further validate the benchmark methodology.
+Appendix G Implementation Details
+G.1 Model Configurations
 Documentation Generation Agent
-We evaluate CodeWiki with two model configurations:
-  • CodeWiki-sonnet-4: Claude Sonnet 4 (claude-sonnet-4-20250514)
+  • Model: Claude Sonnet 4 (claude-sonnet-4-20250514)
   • Temperature: 0.0
-    • Max tokens: 32768
-    • Context window: 200K tokens
-  • CodeWiki-kimi-k2: Kimi K2 Instruct (kimi-k2-instruct-0905)
-    • Temperature: 0.0
-    • Max tokens: 32768
+  • Max tokens: 16384
   • Context window: 200K tokens
@@
   • Binary scoring: 0 (not satisfied) or 1 (satisfied)
-E.2 Processing Parameters
+G.2 Processing Parameters
   • Module decomposition threshold: 32768 tokens per leaf module